### v1

# On the third homology of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$

Abstract.
In this article we study the third homology groups of the special linear group $\mathit{{\rm SL}}_{2}(A)$ and the projective special linear group $\mathit{{\rm PSL}}_{2}(A)$, through the natural surjective map $\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)$. As an application we prove a refined Bloch-Wigner exact sequence over certain local domains.
For any commutative ring $A$ and any non-negative integer $n$, the natural surjective map $\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)$ induces the natural homomorphism of homology groups
$H_{n}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{n}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})$
with 2-power torsion kernel and cokernel. In many problems in algebra and geometry, it is important to understand these kernels and cokernels. In this article we study this map for $n\leq 3$ over certain class of rings. As our first main theorem we prove that if $A$ is a domain, then we have the exact sequence
$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
$\rightarrow H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mu_{2}(A)\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
$\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0,$
where $\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is a quotient of $H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ over a certain 2-torsion subgroup and $\mu_{2}(A)^{\sim}=1$ if ${\rm char}(A)=2$ and $\mu_{2}(A)^{\sim}\simeq\mathbb{Z}/4$ if ${\rm char}(A)\neq 2$ (see Theorem 2.1 for the precise statement). Moreover, we show that the image of the map $\mu_{2}(A)^{\sim}\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ coincides with the image of $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$, where $w={\left(\!\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!0\end{array}\!\!\right)}\in\mathit{{\rm SL}}_{2}(A)$. Observe that the five term exact sequence from the right of the above exact sequence is the well-known five term exact sequence coming from the central extension
$1\rightarrow\mu_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)\rightarrow 1$
As a particular case, if $A$ is a local domain such that its residue field has more than two elements, then we have the exact sequence
$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0$
Furthermore, we will study the above 9-term exact sequence over the ring $\mathbb{Z}[\frac{1}{m}]$, where $m$ is a square free integer (see Proposition 3.1). As the second main result of this article, we will prove a refined Bloch-Wigner exact sequence over certain local rings. More precisely, let $A$ be a local domain such that either $-1\in({A^{\times}})^{2}$ or $|{A^{\times}}/({A^{\times}})^{2}|\leq 4$. Moreover, let $A/\mathfrak{m}_{A}$ be either infinite or it has $p^{d}$ elements where $(p-1)d>6$. Then we prove the refined Bloch-Wigner exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\rightarrow\mathcal{RB}(A)\rightarrow 0,$
where $\mathcal{RB}(A)$ is the refined Bloch group of $A$ defined by Hutchinson [16] (see Theorem 4.4 for the precise statement). In the above two exact sequences the map
$\rho_{\ast}:\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
is induced by the product map
$\rho:\mu_{2}(A)\times\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A),\ \ \ \ (a,X)\mapsto aX.$
If ${\rm char}(A)=2$, clearly this product map is trivial. We will see that when $A$ is a finite field, a quadratically closed field or a real closed field, then the above product map is trivial. If $A$ is a local domain such that its residue field has more than two elements, then the natural product map
$\rho_{\ast}:\mu_{2}(A)\otimes_{\mathbb{Z}}H_{i}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{i+1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
is trivial for $i=0,1$. The following question ask if this is true for $i=2$.
Question 0.1.
Let $A$ be a local domain such that $|A/\mathfrak{m}_{A}|\neq 2$. Is the product map
$\rho_{\ast}:\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
trivial?
We do not know the answer of the above question even over a general infinite field. At the end of Sections 2 and 3, we give examples of non-local domains that this product map is not trivial.
[…]
1. The low dimensional homology of central extensions
Let $\mathcal{A}$ be a central subgroup of a group $\mathcal{G}$. The product map $\rho:\mathcal{A}\times\mathcal{G}\rightarrow\mathcal{G}$, given by $(a,g)\mapsto ag$, induces the natural map $\rho_{\ast}:H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{G},\mathbb{Z})$. By the Künneth formula we have the exact sequence
$\begin{array}[]{c}0\rightarrow\bigoplus_{i=0}^{3}H_{i}(\mathcal{A},\mathbb{Z})\otimes_{\mathbb{Z}}H_{3-i}(\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow 0.\end{array}$
This gives us a natural embedding
${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow\frac{H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})}{\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})\ \oplus H_{2}(\mathcal{A},\mathbb{Z})\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})}.$
It is straightforward to check that the image of the composite
$H_{2}(\mathcal{A},\mathbb{Z})\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\overset{\rho_{\ast}}{\longrightarrow}H_{3}(\mathcal{G},\mathbb{Z})$
is in the image of the composite
$\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\overset{\rho_{\ast}}{\longrightarrow}H_{3}(\mathcal{G},\mathbb{Z})$
(see [29, Proposition 4.4(i), Chap. V]). Thus we get the natural map
${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow H_{3}(\mathcal{G},\mathbb{Z})/\rho_{\ast}(\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})).$
We set
$\widetilde{H}_{3}(\mathcal{G},\mathbb{Z}):=H_{3}(\mathcal{G},\mathbb{Z})/\rho_{\ast}\bigg{(}\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})\oplus{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z})\bigg{)}.$
The following theorem is proved by Eckmann and Hilton in [12], which generalises Ganea’s 6-term exact sequence [29, Theorem 2.2, Chap. V].
Theorem 1.1.
For any central extension $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$, there is a natural map $\tau:\Gamma(\mathcal{A})\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})$ and a natural exact sequence
$H_{4}(\mathcal{Q},\mathbb{Z})\rightarrow\ker(\tau)\rightarrow\widetilde{H}_{3}(\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{Q},\mathbb{Z})\rightarrow{\rm coker}(\tau)\rightarrow H_{2}(\mathcal{G},\mathbb{Z})\rightarrow H_{2}(\mathcal{Q},\mathbb{Z})$
$\rightarrow\mathcal{A}\rightarrow H_{1}(\mathcal{G},\mathbb{Z})\rightarrow H_{1}(\mathcal{Q},\mathbb{Z})\rightarrow 0.$
Proof.
See [12, Theorem 1.1]. ∎
Here $\Gamma$ is Whitehead’s quadratic functor from the category of abelian groups to itself. For more on this functor we refer the reader to [4, §1.2] and [19, §2]. Let us explain the map $\tau$. The inclusion $i:A\hookrightarrow G$ induces the natural map $i_{\ast}:\mathcal{A}\rightarrow\mathcal{G}^{\rm ab}=H_{1}(\mathcal{G},\mathbb{Z})$. By the universal property of $\Gamma$, the quadratic map
$\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z}),\ \ \ a\mapsto a\otimes i_{\ast}(a),$
can be extended to a group homomorphism
$\tau:\Gamma(\mathcal{A})\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z}).$
Since the composite $\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})\overset{\rho_{\ast}}{\longrightarrow}H_{2}(\mathcal{G},\mathbb{Z})$ is trivial, where the second map is given by $a\otimes[g]\mapsto{\boldsymbol{c}}(a,g)$, the composite
[…]
is trivial too. This induces the natural map ${\rm coker}(\tau)\rightarrow H_{2}(\mathcal{G},\mathbb{Z})$. In this article, for calculations purposes, we will use the (left) bar resolution $B_{\bullet}(\mathcal{G})\rightarrow\mathbb{Z}$ of $\mathbb{Z}$ over $\mathbb{Z}[\mathcal{G}]$ to study the integral homology groups $H_{n}(\mathcal{G},\mathbb{Z})$ (see [6, Chap.I, §5]):
$H_{n}(\mathcal{G},\mathbb{Z}):=H_{n}(B_{\bullet}(\mathcal{G})\otimes_{\mathbb{Z}[\mathcal{G}]}\mathbb{Z})=H_{n}(B_{\bullet}(\mathcal{G})_{\mathcal{G}}).$
If any two of the elements $g_{1},\dots,g_{n}\in G$ commute, then we define
${\boldsymbol{c}}(g_{1},\dots,g_{n}):=\overline{\sum_{\sigma\in S_{n}}{\rm sign}(\sigma)[g_{\sigma(1)}|\dots|g_{\sigma(n)}]}\in H_{n}(\mathcal{G},\mathbb{Z}),$
where $S_{n}$ is the symmetric group of degree $n$. For some of properties of this element see [18, Lemma 2.3]. Let $\mathcal{G}$ be a group. Let $K(\mathcal{G},n)$ be an Eilenberg-MacLane space of type $(\mathcal{G},n)$. This is a CW-complex such that
$\pi_{k}(K(\mathcal{G},n),\ast)\simeq\begin{cases}0&\text{if $k\neq n$}\\ \mathcal{G}&\text{if $k=n$}.\end{cases}$
This space is unique up to homotopy. Note that if $n\geq 2$, then $\mathcal{G}$ should be abelian. It is well-known [14, Theorem 21.1] that
$\Gamma(\mathcal{G})\simeq H_{4}(K(\mathcal{G},2),\mathbb{Z}).$
Let $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$ be a central extension. The standard classifying space theory gives a (homotopy theoretic) fibration [9, Chap. 6] of Eilenberg-MacLane spaces $K(\mathcal{A},1)\rightarrow K(\mathcal{G},1)\rightarrow K(\mathcal{Q},1)$ (see [26, 5.1.28]). From this we obtain the fibration
$K(\mathcal{G},1)\rightarrow K(\mathcal{Q},1)\rightarrow K(\mathcal{A},2)$
(see [17, Lemma 3.4.2]). From the Serre spectral sequence of this fibration we obtain the sequence
[…]
This is the sequence
[…]
Let $\mathcal{A}$ be a finite cyclic group. If $2\mid|\mathcal{A}|$, let $\mathcal{A}^{\sim}$ denote the unique non-trivial extension of $\mathcal{A}$ by $\mathbb{Z}/2$. If $2\nmid|\mathcal{A}|$, we define $\mathcal{A}^{\sim}:=\mathcal{A}$. Let $\mathcal{A}$ be an ind-cyclic group, i.e. $\mathcal{A}$ is direct limit of its finite cyclic subgroups. Then by passing to the limit we can define $\mathcal{A}^{\sim}$. Note that if $\mathcal{A}$ has no $2$-torsion, then $\mathcal{A}^{\sim}=\mathcal{A}$. There is always a canonical injective homomorphism $\mathcal{A}\hookrightarrow\mathcal{A}^{\sim}$ whose composition with the projection $\mathcal{A}^{\sim}\rightarrow\mathcal{A}$ coincides with multiplication by $2$.
Lemma 1.2.
If $\mathcal{A}$ is an ind-cyclic group, then $\Gamma(\mathcal{A})\simeq\mathcal{A}^{\sim}$.
Proof.
This follows from the facts that $\Gamma$ commutes with direct limits and $\Gamma(\mathbb{Z}/n)\simeq(\mathbb{Z}/n)^{\sim}$. ∎
Let $\mathcal{A}$ be an abelian group. Let $\sigma_{1}:{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$ be obtained by interchanging $\mathcal{A}$. This map is induced by the involution $\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}$, $a\otimes b\mapsto-b\otimes a$ [20, §2]. Let $\Sigma_{2}^{\prime}=\{1,\sigma^{\prime}\}$ be the symmetric group of order 2. Consider the following action of $\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$:
$(\sigma^{\prime},x)\mapsto-\sigma_{1}(x).$
Proposition 1.3.
For any abelian group $\mathcal{A}$ we have the exact sequence
$\begin{array}[]{c}0\rightarrow\bigwedge_{\mathbb{Z}}^{3}\mathcal{A}\rightarrow H_{3}(\mathcal{A},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})^{\Sigma_{2}^{\prime}}\rightarrow 0.\end{array}$
If $\mathcal{A}$ is an ind-cyclic group, then the action of Σ2superscriptsubscriptnormal-$\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$ is trivial.
Proof.
For the first part see [30, Lemma 5.5] or [5, Section §6]. The second part follows from the first. ∎
Corollary 1.4.
For any integer $m\in\mathbb{Z}$, let $m:\mathcal{A}\rightarrow\mathcal{A}$ be given by $a\mapsto ma$. Then the map $m_{\ast}:H_{3}(\mathcal{A},\mathbb{Z})\rightarrow H_{3}(\mathcal{A},\mathbb{Z})$ induces multiplication by $m^{3}$ on $\bigwedge_{\mathbb{Z}}^{3}\mathcal{A}$ and multiplication by $m^{2}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})^{\Sigma_{2}^{\prime}}$.
Proof.
This follows from the above proposition. ∎
2. Low dimensional homology of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$
We apply Theorem 1.1 to the central extension
[…]
where $\mu_{2}(A)=\{a\in{A^{\times}}:a^{2}=1\}$ and $i:\mu_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)$ is given by $a\mapsto aI_{2}={\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a\end{array}\!\!\right)}$.
Theorem 2.1.
Let $A$ be a ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu_{2}(A)\simeq\mu_{2}(F)$. Then we have the exact sequence
$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
$\rightarrow H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mu_{2}(A)\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
$\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0.$
Moreover, if ${\rm char}(F)\neq 2$, then the image of the maps $\mu_{2}(A)^{\sim}\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ and $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ coincide, where $w:={\left(\!\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!0\end{array}\!\!\right)}$.
Proof.
First we study the homomophism $\tau:\Gamma(\mu_{2}(A))\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$. This map is induced by the quadratic map
$\mu_{2}(A)\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}),\ \ \ \ -1\mapsto(-1)\otimes i_{\ast}(-1),$
where $i_{\ast}:\mu_{2}(A)\simeq H_{1}(\mu_{2}(A),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is induced by $i$. Since $w^{2}=-I_{2}$, we have
$(-1)\otimes i_{\ast}(-1)=(-1)\otimes\overline{[-I_{2}]}=(-1)\otimes\overline{[w^{2}]}=2(-1)\otimes\overline{[w]}=(-1)^{2}\otimes\overline{[w]}=0.$
Thus $\tau$ must be trivial. Hence by Theorem 1.1 we obtain the exact sequence
$H_{4}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\Gamma(\mu_{2}(A))\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})$
$\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow\cdots.$
Therefore to complete the proof of the first part of the theorem, we must prove that the map
$H_{4}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\Gamma(\mu_{2}(A))$
is trivial. Since this map is trivial when $\mu_{2}(A)\simeq\mu_{2}(F)=1$, we may assume that $\mu_{2}(A)\simeq\mu_{2}(F)=\{\pm 1\}\simeq\mathbb{Z}/2$. Let $\overline{F}$ be the algebraic closure of $F$. Then $H_{1}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})=0$ and $H_{2}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$ is isomorphic to $K_{2}^{M}(\overline{F})$, the second Milnor $K$-group of $\overline{F}$ [27, page 286, 2.12]. It is well-known that $K_{2}^{M}(\overline{F})$ is uniquely divisible (see [3, Corollary 1.3]) and thus $\mu_{2}(\overline{F})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})=0$. So we have the exact sequence
$\Gamma(\mu_{2}(\overline{F}))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow 0.$
The surjective map $p:\mathit{{\rm SL}}_{2}(\overline{F})\rightarrow\mathit{{\rm PSL}}_{2}(\overline{F})$ induces the commutative diagram with exact rows
[…]
The first row is the classical Bloch-Wigner exact sequence [10, App. A], [30, Theorem 5.2]. The second exact sequence can be proved as the first exact sequence, where $\widetilde{\mu}(\overline{F}):=\mu(\overline{F})/\mu_{2}(\overline{F})$ (also see Theorem 4.3 below). The left hand side square is obtained from the commutative diagram
[…]
where for a ring $R$, $T(R):=\bigg{\{}{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}|a\in R^{\times}\bigg{\}}$, $\mathit{{\rm PT}}(R):=T(R)/\mu_{2}(R)I_{2}$ and the middle square is induced by the map $\mu(R)\rightarrow T(R)$, $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$. The extension $1\rightarrow\mu_{2}(\overline{F})\rightarrow T(\overline{F})\rightarrow\mathit{{\rm PT}}(\overline{F})\rightarrow 1$ is isomorphic with the extension $1\rightarrow\mu_{2}(\overline{F})\rightarrow\overline{F}^{\times}\overset{()^{2}}{\longrightarrow}\overline{F}^{\times}\rightarrow 1$. The map $\phi=()^{2}:\overline{F}^{\times}\rightarrow\overline{F}^{\times}$ induces the map $\phi_{\ast}:H_{3}(\overline{F}^{\times},\mathbb{Z})\rightarrow H_{3}(\overline{F}^{\times},\mathbb{Z})$. By Corollary 1.4 this leads to multiplication by $2^{2}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))$. This implies that $\ker(\alpha)\simeq\mathbb{Z}/4$ and thus we have the exact sequence
$0\rightarrow\Gamma(\mu_{2}(\overline{F}))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow 0$
Now the first claim follows from the following commutative diagram with exact rows
[…]
For the second claim, first we show that the composite
$H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})$
is trivial. Let denote this map by $\phi_{\ast}$. The group $H_{3}(\langle w\rangle,\mathbb{Z})$ is cyclic of order 4 and is generated by
$\chi:=\overline{[w|w|w]+[w|\!-\!1|w]+[w|\!-\!w|w]}\in H_{3}(B_{\bullet}(\langle w\rangle)_{\langle w\rangle})=H_{3}(\langle w\rangle,\mathbb{Z}),$
(see [32, Proposition 8.1], [25, Proposition 3.25]), where $B_{\bullet}(\langle w\rangle)\rightarrow\mathbb{Z}$ is the bar resolution of $\mathbb{Z}$ over $\langle w\rangle$. We have
$\phi_{\ast}(\chi)=\overline{[w|w|w]+[w|\!-\!w|w]}=\overline{2[w|w|w]}.$
Since $d_{4}([w|w|w|w])={2[w|w|w]}$ in $B_{\bullet}(\mathit{{\rm PSL}}_{2}(A))_{\mathit{{\rm PSL}}_{2}(A)}$, we have $\phi_{\ast}(\chi)=0$. This shows that the above composite is trivial. Now by the commutative diagram
[…]
Corollary 2.2.
Let $A$ be a ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field of ${\rm char}(F)\neq 2$, where $\mu_{2}(A)\simeq\mu_{2}(F)$. Then the map $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is injective.
Proof.
By Theorem 2.1, the map $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is injective. But this map factors through $H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$. Thus we have our claim. ∎
Remark 2.3.
A central extension $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$ is called a stem-extension if the natural map
$\mathcal{A}=H_{1}(\mathcal{A},\mathbb{Z})\rightarrow H_{1}(\mathcal{G},\mathbb{Z})=\mathcal{G}^{\rm ab}$
is trivial [13, §4] and it is called a weak stem-extension if the natural map
$\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})$
is trivial [13, page 103]. The proof of the above theorem shows that if $\mu_{2}(A)=\{\pm 1\}$, then
$1\rightarrow\mu_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)\rightarrow 1$
is a weak stem-extension. But in general it is not a stem-extension (for example take $A=\mathbb{Z}$ or see Proposition 3.1 below).
Corollary 2.4.
Let $A$ be a local ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu_{2}(A)\simeq\mu_{2}(F)$. If $|A/\mathfrak{m}_{A}|\neq 2$, then the sequence
$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0$
is exact.
Proof.
If $|A/\mathfrak{m}_{A}|>3$, then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})=0$ and if $|A/\mathfrak{m}_{A}|=3$, then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq A/\mathfrak{m}_{A}\simeq\mathbb{Z}/3$ [22, Proposition 4.1]. Thus
$\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})=0,\ \ \ \ {{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(A),H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}))=0.$
Hence the claim follows from the above theorem. ∎
Example 2.5.
Let $A$ be a local domain such that $A/\mathfrak{m}_{A}\simeq\mathbb{F}_{2}$. Then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq A/\mathfrak{m}_{A}^{2}$ [22, Proposition 4.1]. By a careful analysis of the composition
$\mu_{2}(A)\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\overset{\simeq}{\longrightarrow}A/\mathfrak{m}_{A}^{2},$
one sees that $-1\in\mu_{2}(A)$ maps to $-\overline{6}\in A/\mathfrak{m}_{A}^{2}$. Since $2\in\mathfrak{m}_{A}$, we have $-\overline{6}=\overline{2}$ in $A/\mathfrak{m}_{A}^{2}$. Thus if $2\in\mathfrak{m}_{A}^{2}$, then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq H_{1}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})$. Otherwise we have the exact sequence
$1\rightarrow\mu_{2}(A)\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0$
Observe that
$\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq\mu_{2}(A)\otimes_{\mathbb{Z}}A/\mathfrak{m}_{A}^{2}\simeq A/(2A+\mathfrak{m}_{A}^{2}).$
If $A=\mathbb{Z}_{(2)}=\{a/b\in\mathbb{Q}:a,b\in\mathbb{Z},2\nmid b\}$ (or $A=\mathbb{Z}_{2}$, the $2$-adic integers), then $2\notin\mathfrak{m}_{A}^{2}$ (by the Nakayama lemma). Thus the sequence
$1\rightarrow\mu_{2}(\mathbb{Z}_{(2)})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow 0$
is exact. Note that $H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq\mathbb{Z}/4$ [22, Example 4.2]. With an argument similar to the proof of Proposition 3.1 below, one can show that $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq H_{2}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})$ and the sequence
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}{\mu_{2}(\mathbb{Z}_{(2)})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0\end{array}$
is exact. We have similar results for $\mathbb{Z}_{2}$. Hence Corollary 2.4 is not necessarily true when $|A/\mathfrak{m}_{A}|=2$.
Example 2.6.
(i) Let $\mathbb{F}_{q}$ be a finite field of odd characteristic. It is well-known that $H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})=0$ when $q\neq 3$ and $H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{3}),\mathbb{Z})\simeq\mathbb{Z}/3$. Moreover, $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})=0$ when $q\neq 9$ and $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{9}),\mathbb{Z})\simeq\mathbb{Z}/3$ [27, page 280]. These facts together with Theorem 2.1 give us the exact sequence
$0\rightarrow\mu_{2}(\mathbb{F}_{q})^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})\rightarrow 0$
For fields of characteristic $2$ the above exact sequence is trivial. Thus this sequence is exact for any finite field.
(ii) If $F$ is quadratically closed, then $H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})$ is isomorphic to $K_{2}^{M}(F)$, the second Milnor $K$-group of $F$. It is well-known that for such fields $K_{2}^{M}(F)$ is uniquely 2-divisible (see [3, Corollary 1.3]). Thus by Corollary 2.4 we obtain the exact sequence
$0\rightarrow\mu_{2}(F)^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\rightarrow 0$
This result first proved by Dupont and Sah in [10, A6, App. A].
(iii) Let $F$ be a real closed field. It is mentioned in [11, page 230] that the sequence
$0\rightarrow\mu_{2}(F)^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\rightarrow 0$
is exact and they refer to the proof of the case of quadratically closed fields discussed in [10, App. A]. We couldn’t really verify this. But in the next section we will give a new proof of this fact (see Corollary 4.7). The same result follows from the results of [7].
Consider the product map
$\rho_{\ast}:\mu_{2}(A)\otimes H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}).$
By the above example, this product map is trivial for finite fields, quadratically closed fields and real closed field. In the introduction we asked if this is true for any local domain such that $A/\mathfrak{m}_{A}\neq\mathbb{F}_{2}$. The following example shows that this map is not necessary trivial over more general domains (see also Example 3.2 below).
Example 2.7.
Let $\mathcal{O}_{-3}$ be the ring of algebraic integers of the imaginary quadratic field $\mathbb{Q}(\sqrt{-3})$. The homology groups of $\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3})$ and $\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3})$ have been calculated in [2, Theorem, page 376] and [28, Theorem 5.7]. We list these homology groups for $1\leq n\leq 3$:
$H_{n}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\simeq\begin{cases}\mathbb{Z}/3&\text{if $n=1$}\\ \mathbb{Z}/4&\text{if $n=2,$}\\ \mathbb{Z}/24\times\mathbb{Z}/6&\text{if $n=3$}\end{cases}$
$H_{n}(\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\simeq\begin{cases}\mathbb{Z}/3&\text{if $n=1$}\\ \mathbb{Z}/4\times\mathbb{Z}/2&\text{if $n=2.$}\\ \mathbb{Z}/6\times\mathbb{Z}/3&\text{if $n=3$}\end{cases}$
From Theorem 2.1 we have the exact sequence
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})}{\mu_{2}(\mathcal{O}_{-3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\rightarrow 0.\end{array}$
From the sizes of the third homology groups of $\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3})$ and $\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3})$ it follows that the product map
$\mu_{2}(\mathcal{O}_{-3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})$
is non-trivial.
3. The low dimensional homology of $\mathit{{\rm SL}}_{2}(\mathbb{Z}[\frac{1}{m}])$ and $\mathit{{\rm PSL}}_{2}(\mathbb{Z}[\frac{1}{m}])$
It is known that $\mathit{{\rm SL}}_{2}(\mathbb{Z})\simeq\mathbb{Z}/4\ast_{\mathbb{Z}/2}\mathbb{Z}/6$ and $\mathit{{\rm PSL}}_{2}(\mathbb{Z})\simeq\mathbb{Z}/2\ast\mathbb{Z}/3$. By the Mayer-Vietoris exact sequence associated to amalgamated product [6, Corollary 7.7, Chap. II] one obtains
$H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})\simeq\mathbb{Z}/12,\ \ \ \ \ H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})=0,$
$H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})\simeq\mathbb{Z}/12,\ \ \ H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}),\mathbb{Z})\simeq\mathbb{Z}/6.$
Using these and Theorem 2.1 we obtain the exact sequence
$0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0$
Proposition 3.1.
Let $m$ be a square free integer and let $A_{m}:=\mathbb{Z}[\frac{1}{m}]$.
(i) If $m$ is even, then $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\simeq H_{1}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})$ and the sequences
$0\rightarrow H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\rightarrow\mu_{2}(A_{m})\rightarrow 0,$
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}{\mu_{2}(A_{m})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\rightarrow 0\end{array}$
are exact.
(ii) If $m$ is odd, then $H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\simeq H_{2}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})$ and the sequences
$0\rightarrow\mu_{2}(A_{m})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\rightarrow 0,$
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}{\mu_{2}(A_{m})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0\end{array}$
are exact.
Proof.
By [22, Proposition 4.3] or [24, Theorem 1.2] we have
$H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\simeq\begin{cases}0&\text{if $2\mid m$, $3\mid m$}\\ \mathbb{Z}/3&\text{if $2\mid m$, $3\nmid m$}\\ \mathbb{Z}/4&\text{if $2\nmid m$, $3\mid m$.}\\ \mathbb{Z}/12&\text{if $2\nmid m$, $3\nmid m$}\end{cases}$
It follows from this that the inclusion $\mathbb{Z}\subseteq A_{m}$ induces the surjective splitting map
$\mathbb{Z}/12\simeq H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})\twoheadrightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z}).$
Now from the commutative diagram
[…]
it follows that the bottom map is trivial. Note that by the exact sequence obtained just before this proposition, the upper map should be trivial. (i) Let $m$ be even. Then from the above calculation of $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$, we see that $\mu_{2}(A_{m})\otimes H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})=0$ and the natural map $\mu_{2}(A_{m})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ is trivial. Thus (i) follows from Theorem 2.1. (ii) Now assume that $m$ is odd. Then $\mu_{2}(A_{m})\otimes H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\simeq\mathbb{Z}/2$. Moreover, from the commutative diagram
[…]
it follows that the bottom map is surjective (note that the upper map is surjective because $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})=0$). Since
$H_{1}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\simeq(\mathbb{Z}/2)^{\alpha}\oplus(\mathbb{Z}/3)^{\beta}$
for some $\alpha$ and $\beta$ (see [31, Corollary 4.4]), from the group structure of $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ one sees that the natural map $\mu_{2}(A_{m})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ should be injective. Hence the claim follows from Theorem 2.1. ∎
Example 3.2.
Let $A_{m}:=\mathbb{Z}[\frac{1}{m}]$, where $m=2$ or $3$.
(i) The product map $\mu_{2}(A_{2})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})$ is trivial: By [23, Corolllary 3.7, Remark 3.8] we have the isomorphism $H_{3}(\mathit{{\rm PSL}}_{2}(A_{2}),\mathbb{Z})\simeq\mathcal{RB}(A_{2})$, where $\mathcal{RB}(A_{2})$ is the refined Bloch group of $A_{2}$ (see the next section). By a result of Coronado and Hutchinson, we have the refined Bloch-Wigner exact sequence
$0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})\rightarrow\mathcal{RB}(A_{2})\rightarrow 0$
(see [8, Proposition 8.31]). From these two results it follows that the desired product map is trivial.
(ii) Now we show that the product map $\mu_{2}(A_{3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})$ is not trivial: We know that $H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/12\oplus\mathbb{Z}/4$. This follows from the cohomology calculation of $\mathit{{\rm SL}}_{2}(A_{3})$ in [1, page 8] and Lemma 3.3 below. Furthermore, by [31, Proposition 4.2], $H^{k}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ is generated by $2$-torsion and $3$-torsion elements for $k\geq 3$. Again by Lemma 3.3 it follows that $H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ is generated by $2$-torsion and $3$-torsion elements. Now let the above product map is trivial. Then from Proposition 3.1 we obtain the exact sequence
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0.\end{array}$
This and the isomorphism $H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/12\oplus\mathbb{Z}/4$ imply that $H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ has an element of order $4$, which is a contradiction.
We learned the following result from Kevin Hutchinson. It seems that this lemma is widely known, but we could not find any suitable reference for it.
Lemma 3.3.
Let $\mathcal{G}$ be a group such that $H_{n}(\mathcal{G},\mathbb{Z})$, $H_{n+1}(\mathcal{G},\mathbb{Z})$ and $H^{n+1}(\mathcal{G},\mathbb{Z})$ are finitely generated. Then
${\rm rank}\ H_{n+1}(\mathcal{G},\mathbb{Z})={\rm rank}\ H^{n+1}(\mathcal{G},\mathbb{Z}),\ \ \ and\ \ \ H_{n}(\mathcal{G},\mathbb{Z})_{\rm tor}\simeq H^{n+1}(\mathcal{G},\mathbb{Z})_{\rm tor}.$
Proof.
This claim easily follows from the Universal Coefficients Theorem
$0\rightarrow{{\rm Ext}_{\mathbb{Z}}^{1}}(H_{n}(\mathcal{G},\mathbb{Z}),\mathbb{Z})\rightarrow H^{n+1}(\mathcal{G},\mathbb{Z})\rightarrow{\rm Hom}(H_{n+1}(\mathcal{G},\mathbb{Z}),\mathbb{Z})\rightarrow 0$
and the facts that for any abelian group $\mathcal{A}$, ${{\rm Ext}_{\mathbb{Z}}^{1}}(\mathbb{Z},\mathcal{A})=0$ and ${{\rm Ext}_{\mathbb{Z}}^{1}}(\mathbb{Z}/n,\mathcal{A})\simeq\mathcal{A}/n\mathcal{A}$. ∎
4. A refined Bloch-Wigner exact sequence
Let $A$ be a commutative ring. For any non-negative integer $n$, let $X_{n}(A^{2})$ be the free abelian group generated by the set of all $(n+1)$-tuples $(\langle{\boldsymbol{v}_{0}}\rangle,\dots,\langle{\boldsymbol{v}_{n}}\rangle)$, where every ${\boldsymbol{v}_{i}}\in A^{2}$ is unimodular and for any two distinct vectors ${\boldsymbol{v}_{i}},{\boldsymbol{v}_{j}}$, the matrix $({\boldsymbol{v}_{i}},{\boldsymbol{v}_{j}})$ is invertible. We consider $X_{l}(A^{2})$ as a left $\mathit{{\rm SL}}_{2}(A)$-module (resp. left $\mathit{{\rm PSL}}_{2}(A)$-module) in a natural way. We convert this action to a right action by the definition $m.g:=g^{-1}m$. The $l$-th differential operator
$\partial_{l}:X_{l}(A^{2})\rightarrow X_{l-1}(A^{2}),\ \ l\geq 1,$
is defined as an alternating sum of face operators which throws away the $i$-th component of generators. Let $\partial_{-1}=\epsilon:X_{0}(A^{2})\rightarrow\mathbb{Z}$ be defined by $\sum_{i}n_{i}(\langle{\boldsymbol{v}_{0,i}}\rangle)\mapsto\sum_{i}n_{i}$. Hence we have the complex
$X_{\bullet}(A^{2})\rightarrow\mathbb{Z}:\ \cdots\longrightarrow X_{2}(A^{2})\overset{\partial_{2}}{\longrightarrow}X_{1}(A^{2})\overset{\partial_{1}}{\longrightarrow}X_{0}(A^{2})\overset{\epsilon}{\longrightarrow}\mathbb{Z}\rightarrow 0.$
Proposition 4.1 (Hutchinson).
If $A$ is a local ring, then $X_{\bullet}(A^{2})\rightarrow\mathbb{Z}$ is exact in dimension $<|A/\mathfrak{m}_{A}|$.
Proof.
See [16, Lemma 3.21]. ∎
From now on we assume that $A$ is a local ring. Let $Z_{2}(A^{2}):=\ker(\partial_{2})$ and set
$\mathcal{RP}(A):=H_{0}(\mathit{{\rm SL}}_{2}(A),Z_{2}(A^{2}))=H_{0}(\mathit{{\rm PSL}}_{2}(A),Z_{2}(A^{2})).$
If the residue field of $A$ has at least 4 elements, then $\mathcal{RP}(A)$ is isomorphic with the quotient of the free $\mathcal{G}_{A}$-module generated by the symbols $[x]$, $x\in\mathcal{W}_{A};=\{a\in A:a(1-a)\in{A^{\times}}\}$, over the subgroup generated by the elements
$[x]-[y]+\langle x\rangle\bigg{[}\frac{y}{x}\bigg{]}-\langle x^{-1}-1\rangle\Bigg{[}\frac{1-x^{-1}}{1-y^{-1}}\Bigg{]}+\langle 1-x\rangle\Bigg{[}\frac{1-x}{1-y}\Bigg{]},\ \ \ \ \ \ x,y,x/y\in\mathcal{W}_{A}$
(see [8, §6] and [21, §2]). The inclusion ${\rm inc}:Z_{2}(A^{2})\rightarrow X_{2}(A^{2})$ induces the map
$\lambda:\mathcal{RP}(A)=Z_{2}(A^{2})_{\mathit{{\rm SL}}_{2}(A)}\overset{\overline{{\rm inc}}}{\longrightarrow}X_{2}(A^{2})_{\mathit{{\rm SL}}_{2}(A)}.$
The orbits of the action of $\mathit{{\rm SL}}_{2}(A)$ on $X_{2}(A)$ is represented by
$\langle a\rangle[\ ]:=(\langle{\boldsymbol{e}_{1}}\rangle,\langle{\boldsymbol{e}_{2}}\rangle,\langle{\boldsymbol{e}_{1}}+a{\boldsymbol{e}_{2}}\rangle),\ \ \ \ \langle a\rangle\in\mathcal{G}_{A}.$
Therefore $X_{2}(A^{2})_{\mathit{{\rm SL}}_{2}(A)}\simeq\mathbb{Z}[\mathcal{G}_{A}]$. The group
$\mathcal{RP}_{1}(A):=\ker\Big{(}\lambda:\mathcal{RP}(A)\rightarrow\mathbb{Z}[\mathcal{G}_{A}]\Big{)}.$
is called the scissors congruence group of $A$. We call
${\rm GW}(A):=H_{0}(\mathit{{\rm SL}}_{2}(A),Z_{1}(A^{2}))$
the Grothendieck-Witt group of $A$ (see [21, Remark 2.2] for an explanation on the choice of this name). Let
$\epsilon:=d_{2,0}^{1}:{\rm GW}(A)\rightarrow\mathbb{Z}.$
The kernel of $\epsilon$ is called the fundamental ideal of $A$ and is denoted by $I(A)$. Let $\mathcal{I}_{A}$ be the augmentation ideal of $\mathbb{Z}[\mathcal{G}_{A}]$. It is easy to see that $\mathcal{I}_{A}$ is generated by the elements $\langle\!\langle a\rangle\!\rangle:=\langle a\rangle-1$, $a\in{A^{\times}}$ and the natural map $\mathbb{Z}[\mathcal{G}_{A}]\rightarrow{\rm GW}(A)$ is surjective and maps $\mathcal{I}_{A}$ onto $I(A)$. If $|A/\mathfrak{m}_{A}|\neq 2$, then
${\rm GW}(A)\simeq\mathbb{Z}[\mathcal{G}_{A}]/\langle\langle\!\langle a\rangle\!\rangle\langle\!\langle 1-a\rangle\!\rangle:a,1-a\in{A^{\times}}\rangle,$
$I(A)\simeq\mathcal{I}_{A}/\langle\langle\!\langle a\rangle\!\rangle\langle\!\langle 1-a\rangle\!\rangle:a,1-a\in{A^{\times}}\rangle$
(see [21, §2]). Let $L_{\bullet}(A^{2})$ be the complex $0\rightarrow Z_{1}(A^{2})\overset{{\rm inc}}{\rightarrow}X_{1}(A^{2})\overset{\partial_{1}}{\rightarrow}X_{0}(A^{2})\rightarrow 0$. From this we obtain the first quadrant spectral sequences
$E^{1}_{p.q}=H_{q}(\mathit{{\rm PSL}}_{2}(A),L_{p}(A^{2}))\Longrightarrow H_{p+q}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z}),$
${E^{\prime}}^{1}_{p.q}=H_{q}(\mathit{{\rm SL}}_{2}(A),L_{p}(A^{2}))\Longrightarrow H_{p+q}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}).$
These spectral sequences have been studies extensively in [23] and [21]. One can show that
$E_{0,q}^{1}\simeq H_{q}(\mathit{{\rm PB}}(A),\mathbb{Z}),\ \ \ \ \ \ E_{1,q}^{1}\simeq H_{q}(\mathit{{\rm PT}}(A),\mathbb{Z}),$
${E^{\prime}}_{0,q}^{1}\simeq H_{q}(B(A),\mathbb{Z}),\ \ \ \ \ \ {E^{\prime}}_{1,q}^{1}\simeq H_{q}(T(A),\mathbb{Z}),$
where
$T(A):=\bigg{\{}\begin{pmatrix}a&0\\ 0&a^{-1}\end{pmatrix}:a\in{A^{\times}}\bigg{\}}\ \ \subseteq\ \ B(A):=\bigg{\{}\begin{pmatrix}a&b\\ 0&a^{-1}\end{pmatrix}:a\in{A^{\times}},b\in A\bigg{\}}.$
and
$\mathit{{\rm PB}}(A):=B(A)/\mu_{2}(A)I_{2},\ \ \ \ \ \ \mathit{{\rm PT}}(A):=T(A)/\mu_{2}(A)I_{2}.$
We denote the matrix $\begin{pmatrix}a&0\\ 0&a^{-1}\end{pmatrix}$ by $D(a)$.
Proposition 4.2 (Hutchinson).
Let $A$ be a local ring (local domain) such that either $A/\mathfrak{m}_{A}$ is infinite or if $|A/\mathfrak{m}_{A}|=p^{d}$, we have $(p-1)d>2(n+1)$ ($(p-1)d>2n$). Then $H_{n}(T(A),\mathbb{Z})\simeq H_{n}(B(A),\mathbb{Z})$ and $H_{n}(\mathit{{\rm PT}}(A),\mathbb{Z})\simeq H_{n}(\mathit{{\rm PB}}(A),\mathbb{Z})$.
Proof.
See [21, Proposition 3.8] and [23, Lemma 2.8]. ∎
From now on we assume that $A$ is a local ring (local domain) such that $A/\mathfrak{m}_{A}$ either is infinite or if it has $p^{d}$ elements, then $(p-1)d>8$ ($(p-1)d>6$). Then we have the isomorphisms
$H_{n}(T(A),\mathbb{Z})\simeq H_{n}(B(A),\mathbb{Z})$
$H_{n}(\mathit{{\rm PT}}(A),\mathbb{Z})\simeq H_{n}(\mathit{{\rm PB}}(A),\mathbb{Z})$
for $n\leq 3$. It is shown in [23, Lemma 2.1] that $E_{2,1}^{2}\simeq\mathcal{RP}_{1}(A)$ and $E_{0,2}^{2}\simeq\displaystyle\frac{{A^{\times}}\wedge{A^{\times}}}{{A^{\times}}\wedge\mu_{2}(A)}$. The kernel of
$\lambda_{1}:=d_{2,1}^{2}:\mathcal{RP}_{1}(A)\rightarrow\frac{{A^{\times}}\wedge{A^{\times}}}{{A^{\times}}\wedge\mu_{2}(A)}$
is denoted by $\mathcal{RB}(A)$ and is called the refined Bloch group of $A$ (see [23, Lemma 2.1 and page 8]). Moreover, it is shown in [21, page 11] that the sequence
[…]
is exact. By studying the spectral sequence $E_{\bullet,\bullet}^{1}$ we proved the following projective refined Bloch-Wigner exact sequence in [23].
Theorem 4.3.
Let $A$ be a local ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu(A)\rightarrow\mu(F)$ is injective. Assume that $A/\mathfrak{m}_{A}$ either is infinite or if it has $p^{d}$ elements, then $(p-1)d>8$ (normal-((($(p-1)d>6$ when $A$ is a domain)normal-))). If either $-1\in{A^{\times}}^{2}$ or $|\mathcal{G}_{A}|\leq 4$, then we have the exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(A),\widetilde{\mu}(A))\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mathcal{RB}(A)\rightarrow 0,$
where $\widetilde{\mu}(A)=\mu(A)/\mu_{2}(A)$.
Proof.
See [23, Theorem 3.1, Theorem 3.2, Corollary 3.3] and [21, Proposition 3.8]. ∎
Now we have the following generalization of Hutchinson’s refined Bloch-Wigner exact sequence.
Theorem 4.4.
Let $A$ be a local ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu(A)\rightarrow\mu(F)$ is injective. Assume that $A/\mathfrak{m}_{A}$ either is infinite or if it has $p^{d}$ elements, then $(p-1)d>8$ (normal-((($(p-1)d>6$ when $A$ is a domain)normal-))). If either $-1\in{A^{\times}}^{2}$ or $|\mathcal{G}_{A}|\leq 4$, then we have the exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)}\rightarrow\mathcal{RB}(A)\rightarrow 0$
Proof.
The composite
$\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\twoheadrightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\twoheadrightarrow\mathcal{RB}(A)$
is surjective. Let $\mathcal{K}$ be the kernel of this map. By Theorem 4.3 we have the commutative diagram with exact rows
[…]
From this diagram and Corollary 2.4 we obtain the exact sequence
$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\mathcal{K}\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(A),\widetilde{\mu}(A))\rightarrow 0$
This clearly implies that $\mathcal{K}\simeq{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}$. Now we study the product map $\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$. Since $A/\mathfrak{m}_{A}$ has at least $4$ elements, we have $A_{A^{\times}}=H_{0}({A^{\times}},A)=0$, where the action of ${A^{\times}}$ on $A$ is given by $a.x:=a^{2}x$. First note that ${E^{\prime}}_{0,1}^{1}\simeq\mathcal{G}_{A}\oplus A_{A^{\times}}=\mathcal{G}_{A}$. The map ${d^{\prime}}_{2,0}^{2}:I(A)\rightarrow\mathcal{G}_{A}$ is given by $\langle\!\langle a\rangle\!\rangle\mapsto\langle a\rangle$ [8, Lemma 4.7]. Thus ${E^{\prime}}_{2,0}^{\infty}\simeq{E^{\prime}}_{2,0}^{2}\simeq I^{2}(A)$. Now from an easy analysis of the spectral sequence ${E^{\prime}}_{\bullet,\bullet}^{1}$ we obtain the exact sequence
$H_{2}(B(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A)\rightarrow 0.$
By Proposition 4.2, $H_{2}(T(A),\mathbb{Z})\simeq H_{2}(B(A),\mathbb{Z})$. The composite
$\mu_{2}(A)\otimes(T(A)\wedge T(A))\simeq\mu_{2}(A)\otimes H_{2}(T(A),\mathbb{Z})\rightarrow\mu_{2}(A)\otimes H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
$\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
is given by $(-1)\otimes(D(a)\wedge D(b))\mapsto{\boldsymbol{c}}(D(-1),D(a),D(b)).$ If $i^{2}=-1$, then
${\boldsymbol{c}}(D(-1),D(a),D(b)=2{\boldsymbol{c}}(D(i),D(a),D(b))=0.$
Let $|\mathcal{G}_{A}|\leq 4$. If $-1\in{A^{\times}}^{2}$, we proceed as in above. If $-1\notin{A^{\times}}^{2}$, then $\mathcal{G}_{A}=\{1,-1,u,-u\}$ for some $u\in{A^{\times}}$. Thus
${\boldsymbol{c}}(D(-1),D(a),D(b))={\boldsymbol{c}}(D(-1),D(-1),D(u))=0.$
In any case, the above composite is trivial and thus
$\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}=\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)}.$
This completes the proof of the theorem. ∎
Remark 4.5.
We actually can calculate the map
$\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
which appears in the above theorem. Let $A$ be a local domain such that ${\rm char}(A)\neq 2$. Let $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in I^{2}(A)$. Then by some tedious calculation one can show that under the surjective map $H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A)$, the cycle class of
$[a|b]+R_{ab}-R_{a}-R_{b}+R_{1}\in B_{2}(\mathit{{\rm SL}}_{2}(A))_{\mathit{{\rm SL}}_{2}(A)}$
maps to $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in I^{2}(A)$, where for any $z\in{A^{\times}}$,
$R_{z}$ $=[w|g^{-1}_{z}]-[w|h^{-1}_{z}]-[h^{-1}_{z}D(z)^{-1}|wh^{-1}_{z}]+[h_{z^{-1}}|w^{-1}]-[h^{-1}_{z}|D(z)^{-1}]$
$+[D(z)h_{z}|D(z)^{-1}]-[D(z)h_{z}|h^{-1}_{z}]+2[h^{-1}_{z}|h^{-1}_{z}]+[h^{-2}_{z}|h^{-2}_{z}]$
$+[D(2)h^{-1}_{z}D(z)^{-1}|D(2)]-[D(2)|h^{-1}_{z}],$
with $g_{z}=\left(\!\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!z\end{array}\!\!\right)$ and $h_{z}=\left(\!\!\begin{array}[]{cc}\!\!1&\!\!z^{-1}\\ \!\!0&\!\!1\end{array}\!\!\right)$. For another cycle that its class maps to $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in I^{2}(A)$ see [15, §7]. Now $(-1)\otimes\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle$ maps in $H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ to the class of the cycle which is obtained from the shuffle product of $[-1]\in B_{1}(\mu_{2}(A))_{\mu_{2}(A)}$ with the cycle $[a|b]+R_{ab}-R_{a}-R_{b}+R_{1}$.
Proposition 4.6.
If $F$ is a real closed field, then the sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{RB}(F)\rightarrow 0$
is exact.
Proof.
It is well-know that for any real closed field $F$, $I(F)\simeq\mathbb{Z}$ which is generated by $\langle\!\langle-1\rangle\!\rangle$ and $I^{2}(F)=2I(F)\simeq 2\mathbb{Z}$. Consider the exact sequence
[…]
(discussed before Theorem 4.3). Note that
$I(F)\otimes_{\mathbb{Z}}\mu_{2}(F)\simeq I(F)/2I(F)\simeq I(F)/I^{2}(F)\simeq\mathcal{G}_{F}\simeq\mathbb{Z}/2.$
Applying the Snake lemma to the commutative diagram
[…]
we obtain the exact sequence
[…]
We show that ${E^{\prime}}_{1,2}^{2}=0$. Consider the differential
${d^{\prime}}_{2,2}^{1}:H_{2}(\mathit{{\rm SL}}_{2}(F),Z_{1}(F^{2}))\rightarrow H_{2}(T(A),\mathbb{Z})\simeq F^{\times}\wedge F^{\times}=F^{>0}\wedge F^{>0}$
(note that $F=\mu_{2}(F)\oplus F^{>0}$ and $F^{>0}=(F^{>0})^{2}$). For any $a,b\in F^{>0}$, let
$Y_{a,b}:=\overline{([D(a)|D(b)]-[D(b)|D(a]))\otimes\bigg{(}(\langle{\boldsymbol{e}_{1}}\rangle,\langle{\boldsymbol{e}_{2}}\rangle)+(\langle{\boldsymbol{e}_{2}}\rangle,\langle{\boldsymbol{e}_{1}}\rangle)\bigg{)}}\in H_{2}(\mathit{{\rm SL}}_{2}(F),Z_{1}(F^{2})).$
It is straightforward to check that
${d^{\prime}}_{2,2}^{1}(Y_{a,b})=2(a\wedge b).$
Since $F^{>0}=(F^{>0})^{2}$, ${d^{\prime}}_{2,2}^{1}$ is surjective and thus ${E^{\prime}}_{1,2}^{2}=0$. Moreover, it is not difficult to show that ${E^{\prime}}_{0,3}^{2}\simeq(\bigwedge_{\mathbb{Z}}^{3}F^{\times})/2\oplus{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))={{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))$ (see [23, Proposition 2.7] and the paragraph after that). Now by an easy analysis of the spectral sequence ${E^{\prime}}_{\bullet,\bullet}^{1}$, we obtain the exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow{E^{\prime}}^{\infty}_{2,1}\rightarrow 0$
From the commutative diagram with exact rows
[…]
it follows that the middle vertical map has a kernel with at most four elements. But we have seen that the composite map
$\mathbb{Z}/4\simeq H_{3}(\langle w\rangle,\mathbb{Z})\hookrightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{RB}(F)$
is trivial (see the proof of Theorem 2.1). This shows that the kernel of the surjective map $H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\twoheadrightarrow\mathcal{RB}(F)$ has exactly four elements. Hence we are done. ∎
Corollary 4.7.
If $F$ is a real closed field, then the sequence
$0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\rightarrow 0$
is exact. In particular, the product map
$\mu_{2}(F)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})$
is trivial.
Proof.
This follows from Proposition 4.6, Theorem 4.3 and Corollary 2.4. ∎

### v3

@@
-On the third homology of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$
+On the connections between the low dimensional homology groups of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$
 Abstract.
-In this article we study the third homology groups of the special linear group $\mathit{{\rm SL}}_{2}(A)$ and the projective special linear group $\mathit{{\rm PSL}}_{2}(A)$, through the natural surjective map $\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)$.
-As an application we prove a refined Bloch-Wigner exact sequence over certain local domains.
-For any commutative ring $A$ and any non-negative integer $n$, the natural surjective map $\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)$ induces the natural homomorphism of homology groups
+In this article we study the low dimensional homology groups of the special linear group $\mathit{{\rm SL}}_{2}(A)$ and the projective special linear group $\mathit{{\rm PSL}}_{2}(A)$, $A$ a domain, through the natural surjective map $\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)$.
+In particular, we study the connection of the first, the second and the third homology groups of these groups over euclidean domains $\mathbb{Z}[\frac{1}{m}]$, $m$ a square free integer, and local domains.
+Let $A$ be a commutative ring with unit and $\mu_{2}(A)$ be the group of $2$-roots of unity in $A$. By studying the Lyndon/Hochschild-Serre spectral sequence associated to the central extension
+$1\rightarrow\mu_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)\rightarrow 1,$
+and using the homology of cyclic groups [6, pages 58-59] and the universal coefficient theorem [6, Exercise 3, §1, Chap III] it follows that, for any non-negative integer $n$, the natural homomorphism of homology groups
 $H_{n}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{n}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})$
-with 2-power torsion kernel and cokernel.
-In many problems in algebra and geometry, it is important to understand these kernels and cokernels.
-In this article we study this map for $n\leq 3$ over certain class of rings.
-As our first main theorem we prove that if $A$ is a domain, then we have the exact sequence
+has 2-power torsion kernel and cokernel. In this article we study this map for $n\leq 3$.
+As our first main theorem we prove that if $A$ is a domain, then we have the 9-term exact sequence
 $0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
@@
 $\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0,$
-where $\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is a quotient of $H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ over a certain 2-torsion subgroup and $\mu_{2}(A)^{\sim}=1$ if ${\rm char}(A)=2$ and $\mu_{2}(A)^{\sim}\simeq\mathbb{Z}/4$ if ${\rm char}(A)\neq 2$ (see Theorem 2.1 for the precise statement).
-Moreover, we show that the image of the map $\mu_{2}(A)^{\sim}\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ coincides with the image of $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$, where $w={\left(\!\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!0\end{array}\!\!\right)}\in\mathit{{\rm SL}}_{2}(A)$.
-Observe that the five term exact sequence from the right of the above exact sequence is the well-known five term exact sequence coming from the central extension
-$1\rightarrow\mu_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)\rightarrow 1$
-As a particular case, if $A$ is a local domain such that its residue field has more than two elements, then we have the exact sequence
-$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0$
-Furthermore, we will study the above 9-term exact sequence over the ring $\mathbb{Z}[\frac{1}{m}]$, where $m$ is a square free integer (see Proposition 3.1).
-As the second main result of this article, we will prove a refined Bloch-Wigner exact sequence over certain local rings.
-More precisely, let $A$ be a local domain such that either $-1\in({A^{\times}})^{2}$ or $|{A^{\times}}/({A^{\times}})^{2}|\leq 4$.
-Moreover, let $A/\mathfrak{m}_{A}$ be either infinite or it has $p^{d}$ elements where $(p-1)d>6$.
-Then we prove the refined Bloch-Wigner exact sequence
-$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\rightarrow\mathcal{RB}(A)\rightarrow 0,$
-where $\mathcal{RB}(A)$ is the refined Bloch group of $A$ defined by Hutchinson [16] (see Theorem 4.4 for the precise statement). In the above two exact sequences the map
+where $\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is a quotient of $H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ over a certain 2-torsion subgroup and
+$\mu_{2}(A)^{\sim}\simeq\begin{cases}0&\text{ if ${\rm char}(A)=2$}\\ \mathbb{Z}/4&\text{ if ${\rm char}(A)\neq 2$}\end{cases}$
+(see Theorem 3.1 for the precise statement).
+Moreover, if ${\rm char}(A)\neq 2$, we show that the image of the map $\mu_{2}(A)^{\sim}\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ coincides with the image of the map $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$, where $w={\left(\!\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!0\end{array}\!\!\right)}\in\mathit{{\rm SL}}_{2}(A)$.
+Observe that the five term exact sequence from the right of the above exact sequence is the well-known five term exact sequence coming from the above central extension [6, Corollary 6.4, Chap.
+VII].
+As first particular case, we study the above 9-term exact sequence over the euclidean domain $A_{m}:=\mathbb{Z}[\frac{1}{m}]$, where $m$ is a square free integer (Proposition 4.1).
+In particular we show that if $m$ is even, then we have the exact sequence
+$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}{\mu_{2}(A_{m})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\rightarrow 0\end{array}$
+and if $m$ is odd, then the sequence
+$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}{\mu_{2}(A_{m})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0\end{array}$
+is exact.
+As second particular case we study the above 9-term exact sequence over local domains (Proposition 5.1).
+In particular, if the residue field of a local domain $A$ is different than $\mathbb{F}_{2}$, then we obtain the exact sequence
+$0\rightarrow\mathbb{Z}/4\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0$
+We will see that this is not necessarily true if the residue field of the local domain is $\mathbb{F}_{2}$.
+For example we show that for $\mathbb{Z}_{(2)}=\{a/b\in\mathbb{Q}:a,b\in\mathbb{Z},2\nmid b\}$, we have the exact sequence
+$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}{\mu_{2}(\mathbb{Z}_{(2)})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0.\end{array}$
+In all the above exact sequences the map
 $\rho_{\ast}:\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
@@
 $\rho:\mu_{2}(A)\times\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A),\ \ \ \ (a,X)\mapsto aX.$
-If ${\rm char}(A)=2$, clearly this product map is trivial.
-We will see that when $A$ is a finite field, a quadratically closed field or a real closed field, then the above product map is trivial.
+We will show that for $A_{2}=\mathbb{Z}[\frac{1}{2}]$, $\rho_{\ast}$ is trivial and for $A_{3}=\mathbb{Z}[\frac{1}{3}]$ the map $\rho_{\ast}$ is non-trivial (see Example 4.4).
 If $A$ is a local domain such that its residue field has more than two elements, then the natural product map
@@
 trivial?
-We do not know the answer of the above question even over a general infinite field. At the end of Sections 2 and 3, we give examples of non-local domains that this product map is not trivial.
-[…]
+We do not know the answer of the above question even over a general infinite field.
+But we show that $\rho_{\ast}$ is trivial for finite fields, quadratically closed fields and real closed fields (Proposition 5.4).
+Here we outline the organization of the present paper.
+In Section 1, we recall some results from the literature over the low dimensional homology groups of central extensions.
+In Section 2 we introduce the scissors congruence group, the Bloch group and formulate the Bloch-Wigner exact sequence over fields.
+In Section 3 we prove our 9-term exact sequence (Theorem 3.1).
+In Section 4, we study the above $9$-term exact sequence over the euclidean domain $\mathbb{Z}[\frac{1}{m}]$.
+Finally in Section 5 we investigate the above $9$-term exact sequence over local rings.
+Notations. In this paper all rings are commutative, except possibly group rings, and have the unit element $1$. For a commutative ring $A$, $\mu(A)$ denotes the group of roots of unity in $A$, i.e.
+$\mu(A):=\{a\in A:\text{there is $n\in\mathbb{N}$ such that $a^{n}=1$}\},$
+and $\mu_{2}(A):=\{a\in A:a^{2}=1\}$. If $\mathcal{B}\rightarrow\mathcal{A}$ is a homomorphism of abelian groups, by $\mathcal{A}/\mathcal{B}$ we mean ${\rm coker}(\mathcal{B}\rightarrow\mathcal{A})$.
 1. The low dimensional homology of central extensions
@@
 $\begin{array}[]{c}0\rightarrow\bigoplus_{i=0}^{3}H_{i}(\mathcal{A},\mathbb{Z})\otimes_{\mathbb{Z}}H_{3-i}(\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow 0.\end{array}$
-This gives us a natural embedding
-${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow\frac{H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})}{\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})\ \oplus H_{2}(\mathcal{A},\mathbb{Z})\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})}.$
+If
+$\widehat{H}_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z}):=\frac{H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})}{\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})\ \oplus H_{2}(\mathcal{A},\mathbb{Z})\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})},$
+then from the above exact sequence we obtain the exact sequence
+$0\rightarrow H_{3}(\mathcal{A},\mathbb{Z})\oplus H_{3}(\mathcal{G},\mathbb{Z})\rightarrow\widehat{H}_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow 0$
+This exact sequence splits naturally, thus we have a natural splitting
+${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\hookrightarrow\widehat{H}_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z}).$
 It is straightforward to check that the image of the composite
@@
 $\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\overset{\rho_{\ast}}{\longrightarrow}H_{3}(\mathcal{G},\mathbb{Z})$
-(see [29, Proposition 4.4(i), Chap. V]). Thus we get the natural map
-${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow H_{3}(\mathcal{G},\mathbb{Z})/\rho_{\ast}(\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})).$
+(see [26, Proposition 4.4(i), Chap. V]). Thus we get the natural composite map
+${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow\widehat{H}_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{G},\mathbb{Z})/\rho_{\ast}(\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})).$
 We set
 $\widetilde{H}_{3}(\mathcal{G},\mathbb{Z}):=H_{3}(\mathcal{G},\mathbb{Z})/\rho_{\ast}\bigg{(}\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})\oplus{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z})\bigg{)}.$
-The following theorem is proved by Eckmann and Hilton in [12], which generalises Ganea’s 6-term exact sequence [29, Theorem 2.2, Chap. V].
-Theorem 1.1.
-For any central extension $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$, there is a natural map $\tau:\Gamma(\mathcal{A})\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})$ and a natural exact sequence
+In this article, for calculations purposes, we will use the (left) bar resolution $B_{\bullet}(\mathcal{G})\rightarrow\mathbb{Z}$ of $\mathbb{Z}$ over $\mathbb{Z}[\mathcal{G}]$ to study the integral homology groups $H_{n}(\mathcal{G},\mathbb{Z})$ (see [6, Chap.I, §5]):
+$H_{n}(\mathcal{G},\mathbb{Z}):=H_{n}(B_{\bullet}(\mathcal{G})\otimes_{\mathbb{Z}[\mathcal{G}]}\mathbb{Z})=H_{n}(B_{\bullet}(\mathcal{G})_{\mathcal{G}}).$
+If any pair of the elements $g_{1},\dots,g_{n}\in G$ commute, then we define
+${\boldsymbol{c}}(g_{1},\dots,g_{n}):=\overline{\sum_{\sigma\in S_{n}}{\rm sign}(\sigma)[g_{\sigma(1)}|\dots|g_{\sigma(n)}]}\in H_{n}(\mathcal{G},\mathbb{Z}),$
+where $S_{n}$ is the symmetric group of degree $n$. If $\mathcal{A}:=\langle g_{1},\dots,g_{n}\rangle\subseteq\mathcal{G}$, then the Pontryagin product (see [6, Chap. V, §5]) induces a natural map
+$\begin{array}[]{c}\bigwedge_{\mathbb{Z}}^{n}\mathcal{A}\rightarrow H_{n}(\mathcal{A},\mathbb{Z})\rightarrow H_{n}(\mathcal{G},\mathbb{Z}).\end{array}$
+It is straightforward to show that the above composite, maps $g_{1}\wedge\cdots\wedge g_{n}$ to ${\boldsymbol{c}}(g_{1},\dots,g_{n})$.
+A function $\psi:\mathcal{A}\rightarrow\mathcal{B}$ of (additive) abelian groups is called a quadratic map if
+(1) for any $a\in\mathcal{A}$, $\psi(a)=\psi(-a)$,
+(2) the function $\mathcal{A}\times\mathcal{A}\rightarrow\mathcal{B}$ with $(a,b)\mapsto\psi(a+b)-\psi(a)-\psi(b)$ is bilinear.
+For any abelian group $\mathcal{A}$, there is a universal quadratic map
+$\gamma:\mathcal{A}\rightarrow\Gamma(\mathcal{A})$
+such that for any quadratic map $\psi:\mathcal{A}\rightarrow\mathcal{B}$, there is a unique group homomorphism $\Psi:\Gamma(\mathcal{A})\rightarrow\mathcal{B}$ such that $\Psi\circ\gamma=\psi$.
+It is easy to see that $\Gamma$ is a functor from the category of abelian groups to itself.
+This functor is called Whitehead’s quadratic functor.
+In fact if $\mathcal{E}$ is the free abelian group generated by the symbols $w(a)$, $a\in\mathcal{A}$, then we can define $\Gamma(\mathcal{A}):=\mathcal{E}/\mathcal{K}$, where $\mathcal{K}$ is the subgroup generated by the relations
+(a) $w(a)-w(-a)$=0,
+(b) $w(a+b+c)-w(a+b)-w(a+c)-w(b+c)+w(a)+w(b)+w(c)=0$.
+The universal quadratic map $\gamma:\mathcal{A}\rightarrow\Gamma(\mathcal{A})$ is given by $a\mapsto\gamma(a):=w(a)+\mathcal{K}$.
+For more on Whitehead’s quadratic functor we refer the reader to [4, §1.2].
+Now let $\mathcal{A}$ be a central subgroup of $\mathcal{G}$.
+The inclusion $i:\mathcal{A}\rightarrow\mathcal{G}$ induces the natural map $i_{\ast}:\mathcal{A}\rightarrow H_{1}(\mathcal{G},\mathbb{Z})$, $a\mapsto{\boldsymbol{c}}(a)$.
+By the universal property of $\Gamma$, the quadratic map
+$\psi:\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z}),\ \ \ a\mapsto a\otimes{\boldsymbol{c}}(a),$
+can be extended to a group homomorphism
+(1.1) $\tau:\Gamma(\mathcal{A})\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z}).$
+[…]
+[…]
+is trivial too. This induces the natural map
+${\rm coker}(\tau)\rightarrow H_{2}(\mathcal{G},\mathbb{Z}).$
+Let $\mathcal{G}$ be a group. Let $K(\mathcal{G},n)$ be an Eilenberg-MacLane space of type $(\mathcal{G},n)$. This is a CW-complex such that
+$\pi_{k}(K(\mathcal{G},n),\ast)\simeq\begin{cases}1&\text{if $k\neq n$}\\ \mathcal{G}&\text{if $k=n$}.\end{cases}$
+This space is unique up to homotopy. Note that for $n\geq 2$, $\mathcal{G}$ should be abelian. It is well-known [14, Theorem 21.1] that
+$\Gamma(\mathcal{A})\simeq H_{4}(K(\mathcal{A},2),\mathbb{Z}).$
+Let $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$ be a central extension.
+The standard classifying space theory gives a (homotopy theoretic) fibration [9, Chap.
+6] of Eilenberg-MacLane spaces
+$K(\mathcal{A},1)\rightarrow K(\mathcal{G},1)\rightarrow K(\mathcal{Q},1)$
+(see [23, 5.1.28]). From this we obtain the fibration
+$K(\mathcal{G},1)\rightarrow K(\mathcal{Q},1)\rightarrow K(\mathcal{A},2)$
+(see [16, Lemma 3.4.2]). From the Serre spectral sequence of this fibration we obtain the sequence
+[…]
+[…]
+$H_{4}(\mathcal{Q},\mathbb{Z})\rightarrow\ker(\tau).$
+The following theorem is proved by Eckmann and Hilton in [12], which generalises Ganea’s 6-term exact sequence [26, Theorem 2.2, Chap. V].
+Theorem 1.1 (Eckmann-Hilton).
+For any central extension $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$, we have the natural $10$-term exact sequence
 $H_{4}(\mathcal{Q},\mathbb{Z})\rightarrow\ker(\tau)\rightarrow\widetilde{H}_{3}(\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{Q},\mathbb{Z})\rightarrow{\rm coker}(\tau)\rightarrow H_{2}(\mathcal{G},\mathbb{Z})\rightarrow H_{2}(\mathcal{Q},\mathbb{Z})$
@@
 See [12, Theorem 1.1]. ∎
-Here $\Gamma$ is Whitehead’s quadratic functor from the category of abelian groups to itself.
-For more on this functor we refer the reader to [4, §1.2] and [19, §2].
-Let us explain the map $\tau$.
-The inclusion $i:A\hookrightarrow G$ induces the natural map $i_{\ast}:\mathcal{A}\rightarrow\mathcal{G}^{\rm ab}=H_{1}(\mathcal{G},\mathbb{Z})$.
-By the universal property of $\Gamma$, the quadratic map
-$\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z}),\ \ \ a\mapsto a\otimes i_{\ast}(a),$
-can be extended to a group homomorphism
-$\tau:\Gamma(\mathcal{A})\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z}).$
-Since the composite $\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})\overset{\rho_{\ast}}{\longrightarrow}H_{2}(\mathcal{G},\mathbb{Z})$ is trivial, where the second map is given by $a\otimes[g]\mapsto{\boldsymbol{c}}(a,g)$, the composite
-[…]
-is trivial too.
-This induces the natural map ${\rm coker}(\tau)\rightarrow H_{2}(\mathcal{G},\mathbb{Z})$.
-In this article, for calculations purposes, we will use the (left) bar resolution $B_{\bullet}(\mathcal{G})\rightarrow\mathbb{Z}$ of $\mathbb{Z}$ over $\mathbb{Z}[\mathcal{G}]$ to study the integral homology groups $H_{n}(\mathcal{G},\mathbb{Z})$ (see [6, Chap.I, §5]):
-$H_{n}(\mathcal{G},\mathbb{Z}):=H_{n}(B_{\bullet}(\mathcal{G})\otimes_{\mathbb{Z}[\mathcal{G}]}\mathbb{Z})=H_{n}(B_{\bullet}(\mathcal{G})_{\mathcal{G}}).$
-If any two of the elements $g_{1},\dots,g_{n}\in G$ commute, then we define
-${\boldsymbol{c}}(g_{1},\dots,g_{n}):=\overline{\sum_{\sigma\in S_{n}}{\rm sign}(\sigma)[g_{\sigma(1)}|\dots|g_{\sigma(n)}]}\in H_{n}(\mathcal{G},\mathbb{Z}),$
-where $S_{n}$ is the symmetric group of degree $n$.
-For some of properties of this element see [18, Lemma 2.3].
-Let $\mathcal{G}$ be a group.
-Let $K(\mathcal{G},n)$ be an Eilenberg-MacLane space of type $(\mathcal{G},n)$.
-This is a CW-complex such that
-$\pi_{k}(K(\mathcal{G},n),\ast)\simeq\begin{cases}0&\text{if $k\neq n$}\\ \mathcal{G}&\text{if $k=n$}.\end{cases}$
-This space is unique up to homotopy. Note that if $n\geq 2$, then $\mathcal{G}$ should be abelian. It is well-known [14, Theorem 21.1] that
-$\Gamma(\mathcal{G})\simeq H_{4}(K(\mathcal{G},2),\mathbb{Z}).$
-Let $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$ be a central extension.
-The standard classifying space theory gives a (homotopy theoretic) fibration [9, Chap.
-6] of Eilenberg-MacLane spaces $K(\mathcal{A},1)\rightarrow K(\mathcal{G},1)\rightarrow K(\mathcal{Q},1)$ (see [26, 5.1.28]).
-From this we obtain the fibration
-$K(\mathcal{G},1)\rightarrow K(\mathcal{Q},1)\rightarrow K(\mathcal{A},2)$
-(see [17, Lemma 3.4.2]). From the Serre spectral sequence of this fibration we obtain the sequence
-[…]
-This is the sequence
-[…]
 Let $\mathcal{A}$ be a finite cyclic group.
@@
 If $2\nmid|\mathcal{A}|$, we define $\mathcal{A}^{\sim}:=\mathcal{A}$.
+Thus if $n=|\mathcal{A}|$, then
+$\mathcal{A}^{\sim}\simeq\begin{cases}\mathbb{Z}/2n&\text{if $2\mid n$}\\ \mathbb{Z}/n&\text{if $2\nmid n$.}\end{cases}$
 Let $\mathcal{A}$ be an ind-cyclic group, i.e.
@@
 Proof.
-This follows from the facts that $\Gamma$ commutes with direct limits and $\Gamma(\mathbb{Z}/n)\simeq(\mathbb{Z}/n)^{\sim}$. ∎
+This follows from the facts that the functor $\Gamma$ commutes with direct limit and
+$\Gamma(\mathbb{Z}/n)\simeq\mathbb{Z}/{\rm gcd}(n^{2},2n)\simeq(\mathbb{Z}/n)^{\sim},$
+where ${\rm gcd}(n^{2},2n)$ is the greatest common divisor of $n^{2}$ and $2n$ (see [4, Corollary 1.2.9]). ∎
 Let $\mathcal{A}$ be an abelian group.
-Let $\sigma_{1}:{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$ be obtained by interchanging $\mathcal{A}$.
-This map is induced by the involution $\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}$, $a\otimes b\mapsto-b\otimes a$ [20, §2].
-Let $\Sigma_{2}^{\prime}=\{1,\sigma^{\prime}\}$ be the symmetric group of order 2.
-Consider the following action of $\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$:
+Let $\sigma_{1}:{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$ be obtained by interchanging the copies of $\mathcal{A}$.
+This map is induced by the involution $\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}$, $a\otimes b\mapsto-b\otimes a$ [18, §2].
+Let $\Sigma_{2}^{\prime}=\{1,\sigma^{\prime}\}$ be the symmetric group of order 2 and consider the following action of $\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$:
 $(\sigma^{\prime},x)\mapsto-\sigma_{1}(x).$
 Proposition 1.3.
-For any abelian group $\mathcal{A}$ we have the exact sequence
+Let $\mathcal{A}$ be an abelian group and $T_{\mathcal{A}}$ be its torsion subgroup. Then we have the exact sequence
+$\begin{array}[]{c}0\rightarrow\bigwedge_{\mathbb{Z}}^{3}\mathcal{A}\rightarrow H_{3}(\mathcal{A},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})^{\Sigma_{2}^{\prime}}\rightarrow 0.\end{array}$
+If $T_{\mathcal{A}}$ is an ind-cyclic group, then $\Sigma_{2}^{\prime}$ acts trivially on ${{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})$ and the exact sequence
+$\begin{array}[]{c}0\rightarrow\bigwedge_{\mathbb{Z}}^{3}\mathcal{A}\rightarrow H_{3}(\mathcal{A},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})\rightarrow 0\end{array}$
+splits naturally.
+Proof.
+By [27, Lemma 5.5] or [5, Section §6] we have the exact sequence
 $\begin{array}[]{c}0\rightarrow\bigwedge_{\mathbb{Z}}^{3}\mathcal{A}\rightarrow H_{3}(\mathcal{A},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})^{\Sigma_{2}^{\prime}}\rightarrow 0.\end{array}$
-If $\mathcal{A}$ is an ind-cyclic group, then the action of Σ2superscriptsubscriptnormal-$\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$ is trivial.
-Proof.
-For the first part see [30, Lemma 5.5] or [5, Section §6]. The second part follows from the first. ∎
+Since ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})\simeq{{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})$, we obtain the first exact sequence.
+Now let $T_{\mathcal{A}}$ be an ind-cyclic group.
+Since ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathbb{Z}/n,\mathbb{Z}/n)\simeq\mathbb{Z}/n$, the action of $\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathbb{Z}/n,\mathbb{Z}/n)$ is trivial.
+Now by passing to the limit, we see that the action of $\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})$ is trivial.
+For the last part note that since $T_{\mathcal{A}}$ is ind-cyclic, $\bigwedge_{\mathbb{Z}}^{3}T_{\mathcal{A}}=0$.
+Now applying the first part to the inclusion $T_{\mathcal{A}}\hookrightarrow\mathcal{A}$, we obtain the following commutative diagram with exact rows
+[…]
+Now from this diagram we obtain a natural splitting map
+${{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})\simeq H_{3}(T_{\mathcal{A}},\mathbb{Z})\rightarrow H_{3}(\mathcal{A},\mathbb{Z}).$
 Corollary 1.4.
@@
 This follows from the above proposition. ∎
+2. The bloch-wigner exact sequence
+Let $A$ be a commutative ring and set $\mathcal{W}_{A}:=\{a\in A:a(1-a)\in{A^{\times}}\}$.
+The scissors congruence group $\mathcal{P}(A)$ of $A$ is defined as the quotient of the free abelian group generated by symbols $[a]$, $a\in\mathcal{W}_{A}$, by the subgroup generated by the elements
+$[a]-[b]+\bigg{[}\frac{b}{a}\bigg{]}-\bigg{[}\frac{1-a^{-1}}{1-b^{-1}}\bigg{]}+\bigg{[}\frac{1-a}{1-b}\bigg{]},$
+where $a,b,a/b\in\mathcal{W}_{A}$. Let $S_{\mathbb{Z}}^{2}({A^{\times}}):=({A^{\times}}\otimes_{\mathbb{Z}}{A^{\times}})/\langle a\otimes b+b\otimes a:a,b\in{A^{\times}}\rangle$. The map
+$\lambda:\mathcal{P}(A)\rightarrow S_{\mathbb{Z}}^{2}({A^{\times}}),\ \ \ \ [a]\mapsto a\otimes(1-a),$
+is well-defined and its kernel is called the Bloch group of $A$ and is denoted by $\mathcal{B}(A)$.
+For any local ring $A$ there is a natural map of graded rings $K_{\bullet}^{M}(A)\to K_{\bullet}(A)$, from Milnor $K$-theory to $K$-theory. It is easy to see that
+$K_{0}^{M}(A)\simeq K_{0}(A)\simeq\mathbb{Z},\ \ \ \ K_{1}^{M}(A)\simeq K_{1}(A)\simeq{A^{\times}}.$
+Moreover, Matsumoto proved that for a field $F$
+$K_{2}(F)\simeq K_{2}^{M}(F)\simeq({A^{\times}}\otimes_{\mathbb{Z}}{A^{\times}})/\langle a\otimes(1-a):a\in\mathcal{W}_{F}\rangle$
+(see [23, Theorem 4.3.15]).
+The indecomposable $K_{3}^{\rm ind}(F)$ of a field $F$ is defined as follows:
+$K_{3}^{\rm ind}(F):=K_{3}(F)/K_{3}^{M}(F).$
+The Bloch group and the indecomposable part of the third $K$-group are deeply connected.
+Theorem 2.1 (Bloch-Wigner exact sequence).
+For any field $F$, we have a natural exact sequence
+$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))^{\sim}\rightarrow K_{3}^{\rm ind}(F)\rightarrow\mathcal{B}(F)\rightarrow 0$
+Proof.
+The case of infinite fields has been proved in [27, Theorem 5.2] and the case of finite fields has been settled in [15, Corollary 7.5]. ∎
+For any filed $F$, there are natural maps $H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow K_{2}(F)$ and $H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow K_{3}^{\rm ind}(F)$ (see [15, §5]).
+Proposition 2.2.
+If $F$ is a quadratically closed field, then we have natural isomorphisms
+$K_{2}(F)\simeq K_{2}^{M}(F)\simeq H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z}),\ \ \ \ \ K_{3}^{\rm ind}(F)\simeq H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z}).$
+Furthermore, $K_{2}(F)$ is uniquely $2$-divisible.
+Proof.
+See [24, page 286, 2.12], [17, Proposition 6.4] and [3, Proposition 1.2]. ∎
+The following classical Bloch-Wigner exact sequence will be used in the proof of Theorem 3.1.
+Corollary 2.3 (Bloch-Wigner, Dupont-Sah).
+If $F$ is a quadritically closed field, then we have the exact sequence
+${0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{B}(F)\rightarrow 0.}$
+Proof.
+For a direct proof see [10, Appedix A]. See also [27, Theorem 5.2]. ∎
+Note that the map $\mu(F)\rightarrow\mathit{{\rm SL}}_{2}(F)$ given by $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$ induces the natural map
+$H_{3}(\mu(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z}).$
+By Proposition 1.3, $H_{3}(\mu(F),\mathbb{Z})\simeq{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))$. Now the left map in the above classical Bloch-Wigner exact sequence is the map
+${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))\simeq H_{3}(\mu(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z}).$
+Proposition 2.4 (Coronado [7]).
+Let $F$ be a real closed field. Then we have the Bloch-Wigner exact sequence
+${0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{B}(F)\rightarrow 0.}$
+In particular, $H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\simeq K_{3}^{\rm ind}(F)$
+Proof.
+See [7, Theorem 6.1] and its proof. ∎
 3. Low dimensional homology of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$
@@
 […]
-where $\mu_{2}(A)=\{a\in{A^{\times}}:a^{2}=1\}$ and $i:\mu_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)$ is given by $a\mapsto aI_{2}={\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a\end{array}\!\!\right)}$.
+where $\mu_{2}(A)=\{a\in{A^{\times}}:a^{2}=1\}$ and the map $i$ is given by $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a\end{array}\!\!\right)}={\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$.
+Let
+$\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}):=\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\rho_{\ast}\Big{(}\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\oplus{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(A),H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}))\Big{)}},$
+where $\rho:\mu_{2}(A)\times\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)$ is the product map and is given by $(a,X)\mapsto aX$. We recall that for a field $F$,
+$\mu_{2}(F)^{\sim}\simeq\begin{cases}0&\text{if ${\rm char}(F)=2$}\\ \mathbb{Z}/4&\text{if ${\rm char}(F)\neq 2$}.\end{cases}$
+Now we are ready to state and prove our first main result.
 Theorem 3.1.
-Let $A$ be a ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu_{2}(A)\simeq\mu_{2}(F)$. Then we have the exact sequence
+Let $A$ be a ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu_{2}(A)\simeq\mu_{2}(F)$. Then we have the $9$-term exact sequence
 $0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
@@
 Proof.
-First we study the homomophism $\tau:\Gamma(\mu_{2}(A))\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$. This map is induced by the quadratic map
+First we study the homomophism $\tau:\Gamma(\mu_{2}(A))\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ (see (1.1)).
+Note that by Lemma 1.2, $\Gamma(\mu_{2}(A))\simeq\mu_{2}(A)^{\sim}$.
+The map $\tau$ is induced by the quadratic map
 $\mu_{2}(A)\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}),\ \ \ \ -1\mapsto(-1)\otimes i_{\ast}(-1),$
-where $i_{\ast}:\mu_{2}(A)\simeq H_{1}(\mu_{2}(A),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is induced by $i$. Since $w^{2}=-I_{2}$, we have
-$(-1)\otimes i_{\ast}(-1)=(-1)\otimes\overline{[-I_{2}]}=(-1)\otimes\overline{[w^{2}]}=2(-1)\otimes\overline{[w]}=(-1)^{2}\otimes\overline{[w]}=0.$
-Thus $\tau$ must be trivial. Hence by Theorem 1.1 we obtain the exact sequence
+where $i_{\ast}:\mu_{2}(A)\simeq H_{1}(\mu_{2}(A),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is given by $-1\mapsto{\boldsymbol{c}}(-I_{2})$. Since $w^{2}=-I_{2}$, we have
+$(-1)\otimes i_{\ast}(-1)=(-1)\otimes{\boldsymbol{c}}(-I_{2})=(-1)\otimes{\boldsymbol{c}}(w^{2})=(-1)\otimes 2{\boldsymbol{c}}(w)=(-1)^{2}\otimes{\boldsymbol{c}}(w)=0.$
+Thus $\tau$ must be trivial. Now by Theorem 1.1 we obtain the exact sequence
 $H_{4}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\Gamma(\mu_{2}(A))\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})$
@@
 Therefore to complete the proof of the first part of the theorem, we must prove that the map
-$H_{4}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\Gamma(\mu_{2}(A))$
-is trivial.
-Since this map is trivial when $\mu_{2}(A)\simeq\mu_{2}(F)=1$, we may assume that $\mu_{2}(A)\simeq\mu_{2}(F)=\{\pm 1\}\simeq\mathbb{Z}/2$.
+$H_{4}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\Gamma(\mu_{2}(A))\simeq\mu_{2}(A)^{\sim}$
+is trivial. Since this map is trivial when $\mu_{2}(A)\simeq\mu_{2}(F)=1$, we may assume that $\mu_{2}(A)\simeq\mu_{2}(F)=\{\pm 1\}\simeq\mathbb{Z}/2$.
 Let $\overline{F}$ be the algebraic closure of $F$.
-Then $H_{1}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})=0$ and $H_{2}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$ is isomorphic to $K_{2}^{M}(\overline{F})$, the second Milnor $K$-group of $\overline{F}$ [27, page 286, 2.12].
-It is well-known that $K_{2}^{M}(\overline{F})$ is uniquely divisible (see [3, Corollary 1.3]) and thus $\mu_{2}(\overline{F})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})=0$.
-So we have the exact sequence
-$\Gamma(\mu_{2}(\overline{F}))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow 0.$
+Then $H_{1}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})=0$ and $H_{2}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$ is uniquely 2-divisible by Proposition 2.2.
+These imply that ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(\overline{F}),H_{1}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z}))$ and $\mu_{2}(\overline{F})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$ are trivial.
+Hence by Theorem 1.1, we have the exact sequence
+(3.1) $\Gamma(\mu_{2}(\overline{F}))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow 0.$
+Observe that $\Gamma(\mu_{2}(\overline{F}))\simeq\mu_{2}(\overline{F})^{\sim}\simeq\mathbb{Z}/4$ (Lemma 1.2).
 The surjective map $p:\mathit{{\rm SL}}_{2}(\overline{F})\rightarrow\mathit{{\rm PSL}}_{2}(\overline{F})$ induces the commutative diagram with exact rows
+00{0}${{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))}$${H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})}$${\mathcal{\mathcal{B}}(\overline{F})}$00{0}00{0}${{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(\overline{F}),\widetilde{\mu}(\overline{F}))}$${H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})}$${\mathcal{B}(\overline{F})}$${0.}$$\scriptstyle{\alpha}$$\scriptstyle{p_{\ast}}$
+The first row is the classical Bloch-Wigner exact sequence discussed in the previous section (Corollary 2.3).
+The second exact sequence can be proved as the first exact sequence, where $\widetilde{\mu}(\overline{F}):=\mu(\overline{F})/\mu_{2}(\overline{F})$ (see [10, Theorem 4.10 and Appendix A].
+Also see [20, Corollary 3.5] below).
+The left hand side square is obtained from the commutative diagram
+${{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))}$${H_{3}(\mu(\overline{F}),\mathbb{Z})}$${H_{3}({\rm T}(\overline{F}),\mathbb{Z})}$${H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})}$${{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(\overline{F}),\widetilde{\mu}(\overline{F}))}$${H_{3}(\widetilde{\mu}(\overline{F}),\mathbb{Z})}$${H_{3}(\mathit{{\rm PT}}(\overline{F}),\mathbb{Z})}$${H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z}),}$$\scriptstyle{\alpha}$$\scriptstyle{p_{\ast}}$$\scriptstyle{p_{\ast}}$$\scriptstyle{p_{\ast}}$
+where for a ring $R$,
+${\rm T}(R):=\bigg{\{}{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}|a\in R^{\times}\bigg{\}},\ \ \ \ \ \mathit{{\rm PT}}(R):={\rm T}(R)/\mu_{2}(R)I_{2}$
+and the middle square is induced by the map $\mu(R)\rightarrow{\rm T}(R)$, $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$. The extension
+$1\rightarrow\mu_{2}(\overline{F})\rightarrow{\rm T}(\overline{F})\rightarrow\mathit{{\rm PT}}(\overline{F})\rightarrow 1$
+is isomorphic with the extension
+$1\rightarrow\mu_{2}(\overline{F})\rightarrow\overline{F}^{\times}\overset{()^{2}}{\longrightarrow}\overline{F}^{\times}\rightarrow 1$
+The map $\phi=()^{2}:\overline{F}^{\times}\rightarrow\overline{F}^{\times}$ induces the map $\phi_{\ast}:H_{3}(\overline{F}^{\times},\mathbb{Z})\rightarrow H_{3}(\overline{F}^{\times},\mathbb{Z})$.
+By Proposition 1.3 we have the split exact sequence
+$\begin{array}[]{c}0\rightarrow\bigwedge_{\mathbb{Z}}^{3}\overline{F}^{\times}\rightarrow H_{3}(\overline{F}^{\times},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))\rightarrow 0.\end{array}$
+By Corollary 1.4 this leads to multiplication by $2^{2}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))$. Since
+${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))\simeq\mu(\overline{F}),$
+the above fact implies that $\ker(\alpha)\simeq\mathbb{Z}/4$.
+On the other hand by the Snake lemma $\ker(\alpha)\simeq\ker(p_{\ast})$.
+Thus the left map of the exact sequence (3.1) must be injective and so we have the exact sequence
+$0\rightarrow\Gamma(\mu_{2}(\overline{F}))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow 0$
+Observe that
+$\Gamma(\mu_{2}(\overline{F}))\simeq\mu_{2}(\overline{F})^{\sim}\hookrightarrow\mu(\overline{F})^{\sim}\simeq\mu(\overline{F})\simeq{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F})).$
+The composite $A\rightarrow F\hookrightarrow\overline{F}$ induces the morphism of extensions
 […]
-The first row is the classical Bloch-Wigner exact sequence [10, App.
-A], [30, Theorem 5.2].
-The second exact sequence can be proved as the first exact sequence, where $\widetilde{\mu}(\overline{F}):=\mu(\overline{F})/\mu_{2}(\overline{F})$ (also see Theorem 4.3 below).
-The left hand side square is obtained from the commutative diagram
-[…]
-where for a ring $R$, $T(R):=\bigg{\{}{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}|a\in R^{\times}\bigg{\}}$, $\mathit{{\rm PT}}(R):=T(R)/\mu_{2}(R)I_{2}$ and the middle square is induced by the map $\mu(R)\rightarrow T(R)$, $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$.
-The extension $1\rightarrow\mu_{2}(\overline{F})\rightarrow T(\overline{F})\rightarrow\mathit{{\rm PT}}(\overline{F})\rightarrow 1$ is isomorphic with the extension $1\rightarrow\mu_{2}(\overline{F})\rightarrow\overline{F}^{\times}\overset{()^{2}}{\longrightarrow}\overline{F}^{\times}\rightarrow 1$.
-The map $\phi=()^{2}:\overline{F}^{\times}\rightarrow\overline{F}^{\times}$ induces the map $\phi_{\ast}:H_{3}(\overline{F}^{\times},\mathbb{Z})\rightarrow H_{3}(\overline{F}^{\times},\mathbb{Z})$.
-By Corollary 1.4 this leads to multiplication by $2^{2}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))$.
-This implies that $\ker(\alpha)\simeq\mathbb{Z}/4$ and thus we have the exact sequence
-$0\rightarrow\Gamma(\mu_{2}(\overline{F}))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow 0$
-Now the first claim follows from the following commutative diagram with exact rows
-[…]
+Now by Theorem 1.1 we obtain the commutative diagram with exact rows
+${H_{4}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})}$${\Gamma(\mu_{2}(A))}$${\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}$${H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})}$00{0}${\Gamma(\mu_{2}(\overline{F}))}$${H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})}$${H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})}$${0,}$
+where the exactness of the second row has just been shown. Since $\mu_{2}(A)\simeq\mu_{2}(F)=\mu_{2}(\overline{F})$, we have
+$\Gamma(\mu_{2}(A))\simeq\Gamma(\mu_{2}(\overline{F})).$
+Now from the above diagram it follows that the map $\Gamma(\mu_{2}(A))\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is injective. This complete the proof of the first claim.
 For the second claim, first we show that the composite
 $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})$
-is trivial. Let denote this map by $\phi_{\ast}$. The group $H_{3}(\langle w\rangle,\mathbb{Z})$ is cyclic of order 4 and is generated by
+is trivial. Let denote this map by $\varphi_{\ast}$. The group $H_{3}(\langle w\rangle,\mathbb{Z})$ is cyclic of order 4 and is generated by
 $\chi:=\overline{[w|w|w]+[w|\!-\!1|w]+[w|\!-\!w|w]}\in H_{3}(B_{\bullet}(\langle w\rangle)_{\langle w\rangle})=H_{3}(\langle w\rangle,\mathbb{Z}),$
-(see [32, Proposition 8.1], [25, Proposition 3.25]), where $B_{\bullet}(\langle w\rangle)\rightarrow\mathbb{Z}$ is the bar resolution of $\mathbb{Z}$ over $\langle w\rangle$. We have
-$\phi_{\ast}(\chi)=\overline{[w|w|w]+[w|\!-\!w|w]}=\overline{2[w|w|w]}.$
-Since $d_{4}([w|w|w|w])={2[w|w|w]}$ in $B_{\bullet}(\mathit{{\rm PSL}}_{2}(A))_{\mathit{{\rm PSL}}_{2}(A)}$, we have $\phi_{\ast}(\chi)=0$.
-This shows that the above composite is trivial.
-Now by the commutative diagram
-[…]
-[…]
+(see [29, Proposition 8.1] or [22, Proposition 3.25]), where $B_{\bullet}(\langle w\rangle)\rightarrow\mathbb{Z}$ is the bar resolution of $\mathbb{Z}$ over $\langle w\rangle$. We have
+$\varphi_{\ast}(\chi)=\overline{[w|w|w]+[w|\!-\!w|w]}=\overline{2[w|w|w]}.$
+Since
+$d_{4}([w|w|w|w])={2[w|w|w]}$
+in $B_{\bullet}(\mathit{{\rm PSL}}_{2}(A))_{\mathit{{\rm PSL}}_{2}(A)}$, we have $\varphi_{\ast}(\chi)=0$. This shows that the above composite is trivial. Now by the commutative diagram
+${H_{3}(\langle w\rangle,\mathbb{Z})}$${\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}$${H_{3}(\langle w\rangle,\mathbb{Z})}$${H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z}),}$
+it is sufficient to prove that the map $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$ is injective.
+Let $\mathfrak{i}^{2}=-1$.
+Then $w$ is conjugate to ${\left(\!\!\begin{array}[]{cc}\!\!\mathfrak{i}&\!\!0\\ \!\!0&\!\!\mathfrak{i}^{-1}\end{array}\!\!\right)}$.
+It is known by the classical Bloch-Wigner exact sequence (Corollary 2.3) that the map
+$H_{3}(\mu(\overline{F}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$
+which is induced by $\mu(\overline{F})\rightarrow\mathit{{\rm SL}}_{2}(\overline{F})$, $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$, is injective.
+Thus the map
+$H_{3}(\langle\mathfrak{i}\rangle,\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$
+is injective. This proves our claim. ∎
+Remark 3.2.
+The above theorem is also valid if we replace $\mathit{{\rm SL}}_{2}(A)$ with the elementary subgroup ${\rm E}_{2}(A)$.
+The proof stays the same.
+Just one should note that over a field $F$, ${\rm E}_{2}(F)=\mathit{{\rm SL}}_{2}(F)$.
 Corollary 3.3.
@@
 ∎
-Remark 2.3.
-A central extension $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$ is called a stem-extension if the natural map
-$\mathcal{A}=H_{1}(\mathcal{A},\mathbb{Z})\rightarrow H_{1}(\mathcal{G},\mathbb{Z})=\mathcal{G}^{\rm ab}$
-is trivial [13, §4] and it is called a weak stem-extension if the natural map
-$\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})$
-is trivial [13, page 103]. The proof of the above theorem shows that if $\mu_{2}(A)=\{\pm 1\}$, then
-$1\rightarrow\mu_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)\rightarrow 1$
-is a weak stem-extension. But in general it is not a stem-extension (for example take $A=\mathbb{Z}$ or see Proposition 3.1 below).
-Corollary 2.4.
-Let $A$ be a local ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu_{2}(A)\simeq\mu_{2}(F)$. If $|A/\mathfrak{m}_{A}|\neq 2$, then the sequence
-$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0$
-is exact.
-Proof.
-If $|A/\mathfrak{m}_{A}|>3$, then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})=0$ and if $|A/\mathfrak{m}_{A}|=3$, then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq A/\mathfrak{m}_{A}\simeq\mathbb{Z}/3$ [22, Proposition 4.1].
-Thus
-$\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})=0,\ \ \ \ {{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(A),H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}))=0.$
-Hence the claim follows from the above theorem. ∎
-Example 2.5.
-Let $A$ be a local domain such that $A/\mathfrak{m}_{A}\simeq\mathbb{F}_{2}$.
-Then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq A/\mathfrak{m}_{A}^{2}$ [22, Proposition 4.1].
-By a careful analysis of the composition
-$\mu_{2}(A)\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\overset{\simeq}{\longrightarrow}A/\mathfrak{m}_{A}^{2},$
-one sees that $-1\in\mu_{2}(A)$ maps to $-\overline{6}\in A/\mathfrak{m}_{A}^{2}$.
-Since $2\in\mathfrak{m}_{A}$, we have $-\overline{6}=\overline{2}$ in $A/\mathfrak{m}_{A}^{2}$.
-Thus if $2\in\mathfrak{m}_{A}^{2}$, then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq H_{1}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})$.
-Otherwise we have the exact sequence
-$1\rightarrow\mu_{2}(A)\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0$
-Observe that
-$\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq\mu_{2}(A)\otimes_{\mathbb{Z}}A/\mathfrak{m}_{A}^{2}\simeq A/(2A+\mathfrak{m}_{A}^{2}).$
-If $A=\mathbb{Z}_{(2)}=\{a/b\in\mathbb{Q}:a,b\in\mathbb{Z},2\nmid b\}$ (or $A=\mathbb{Z}_{2}$, the $2$-adic integers), then $2\notin\mathfrak{m}_{A}^{2}$ (by the Nakayama lemma). Thus the sequence
-$1\rightarrow\mu_{2}(\mathbb{Z}_{(2)})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow 0$
-is exact.
-Note that $H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq\mathbb{Z}/4$ [22, Example 4.2].
-With an argument similar to the proof of Proposition 3.1 below, one can show that $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq H_{2}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})$ and the sequence
-$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}{\mu_{2}(\mathbb{Z}_{(2)})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0\end{array}$
-is exact. We have similar results for $\mathbb{Z}_{2}$. Hence Corollary 2.4 is not necessarily true when $|A/\mathfrak{m}_{A}|=2$.
-Example 2.6.
-(i) Let $\mathbb{F}_{q}$ be a finite field of odd characteristic.
-It is well-known that $H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})=0$ when $q\neq 3$ and $H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{3}),\mathbb{Z})\simeq\mathbb{Z}/3$.
-Moreover, $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})=0$ when $q\neq 9$ and $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{9}),\mathbb{Z})\simeq\mathbb{Z}/3$ [27, page 280].
-These facts together with Theorem 2.1 give us the exact sequence
-$0\rightarrow\mu_{2}(\mathbb{F}_{q})^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})\rightarrow 0$
-For fields of characteristic $2$ the above exact sequence is trivial. Thus this sequence is exact for any finite field.
-(ii) If $F$ is quadratically closed, then $H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})$ is isomorphic to $K_{2}^{M}(F)$, the second Milnor $K$-group of $F$.
-It is well-known that for such fields $K_{2}^{M}(F)$ is uniquely 2-divisible (see [3, Corollary 1.3]).
-Thus by Corollary 2.4 we obtain the exact sequence
-$0\rightarrow\mu_{2}(F)^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\rightarrow 0$
-This result first proved by Dupont and Sah in [10, A6, App. A].
-(iii) Let $F$ be a real closed field. It is mentioned in [11, page 230] that the sequence
-$0\rightarrow\mu_{2}(F)^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\rightarrow 0$
-is exact and they refer to the proof of the case of quadratically closed fields discussed in [10, App.
-A].
-We couldn’t really verify this.
-But in the next section we will give a new proof of this fact (see Corollary 4.7).
-The same result follows from the results of [7].
-Consider the product map
-$\rho_{\ast}:\mu_{2}(A)\otimes H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}).$
-By the above example, this product map is trivial for finite fields, quadratically closed fields and real closed field.
-In the introduction we asked if this is true for any local domain such that $A/\mathfrak{m}_{A}\neq\mathbb{F}_{2}$.
-The following example shows that this map is not necessary trivial over more general domains (see also Example 3.2 below).
-Example 2.7.
-Let $\mathcal{O}_{-3}$ be the ring of algebraic integers of the imaginary quadratic field $\mathbb{Q}(\sqrt{-3})$.
-The homology groups of $\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3})$ and $\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3})$ have been calculated in [2, Theorem, page 376] and [28, Theorem 5.7].
-We list these homology groups for $1\leq n\leq 3$:
-$H_{n}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\simeq\begin{cases}\mathbb{Z}/3&\text{if $n=1$}\\ \mathbb{Z}/4&\text{if $n=2,$}\\ \mathbb{Z}/24\times\mathbb{Z}/6&\text{if $n=3$}\end{cases}$
-$H_{n}(\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\simeq\begin{cases}\mathbb{Z}/3&\text{if $n=1$}\\ \mathbb{Z}/4\times\mathbb{Z}/2&\text{if $n=2.$}\\ \mathbb{Z}/6\times\mathbb{Z}/3&\text{if $n=3$}\end{cases}$
-From Theorem 2.1 we have the exact sequence
-$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})}{\mu_{2}(\mathcal{O}_{-3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\rightarrow 0.\end{array}$
-From the sizes of the third homology groups of $\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3})$ and $\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3})$ it follows that the product map
-$\mu_{2}(\mathcal{O}_{-3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})$
-is non-trivial.
 4. The low dimensional homology of $\mathit{{\rm SL}}_{2}(\mathbb{Z}[\frac{1}{m}])$ and $\mathit{{\rm PSL}}_{2}(\mathbb{Z}[\frac{1}{m}])$
 It is known that $\mathit{{\rm SL}}_{2}(\mathbb{Z})\simeq\mathbb{Z}/4\ast_{\mathbb{Z}/2}\mathbb{Z}/6$ and $\mathit{{\rm PSL}}_{2}(\mathbb{Z})\simeq\mathbb{Z}/2\ast\mathbb{Z}/3$.
-By the Mayer-Vietoris exact sequence associated to amalgamated product [6, Corollary 7.7, Chap.
+By the Mayer-Vietoris exact sequence associated to a amalgamated product [6, Corollary 7.7, Chap.
 II] one obtains
@@
 Using these and Theorem 3.1 we obtain the exact sequence
-$0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0$
+(4.1) $0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0,$
+which will be used in the proof of the following result.
 Proposition 4.1.
@@
 Proof.
-By [22, Proposition 4.3] or [24, Theorem 1.2] we have
+By [19, Proposition 4.3] or [21, Theorem 1.2] we have
 $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\simeq\begin{cases}0&\text{if $2\mid m$, $3\mid m$}\\ \mathbb{Z}/3&\text{if $2\mid m$, $3\nmid m$}\\ \mathbb{Z}/4&\text{if $2\nmid m$, $3\mid m$.}\\ \mathbb{Z}/12&\text{if $2\nmid m$, $3\nmid m$}\end{cases}$
@@
 Now from the commutative diagram
-[…]
+${{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(\mathbb{Z}),H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z}))}$${H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})}$${{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(A_{m}),H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z}))}$${\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}{\mu_{2}(A_{m})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}}$$\scriptstyle{\rho_{\ast}=0}$$\scriptstyle{\rho_{\ast}}$
 it follows that the bottom map is trivial.
-Note that by the exact sequence obtained just before this proposition, the upper map should be trivial.
-(i) Let $m$ be even.
-Then from the above calculation of $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$, we see that $\mu_{2}(A_{m})\otimes H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})=0$ and the natural map $\mu_{2}(A_{m})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ is trivial.
-Thus (i) follows from Theorem 2.1.
-(ii) Now assume that $m$ is odd.
-Then $\mu_{2}(A_{m})\otimes H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\simeq\mathbb{Z}/2$.
-Moreover, from the commutative diagram
-[…]
-it follows that the bottom map is surjective (note that the upper map is surjective because $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})=0$). Since
+In the above diagram, first note that since $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})=0$, we have $H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})=\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})}{\mu_{2}(\mathbb{Z})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})}$ and thus the above diagram is commutative.
+Second, by the exact sequence (4.1), the upper map in the above diagram should be trivial.
+(i) Let $m$ be even. Then from the above calculation of $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$, we see that
+$\mu_{2}(A_{m})\otimes H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})=0$
+and the natural map $\mu_{2}(A_{m})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ is trivial. Thus the claim follows from Theorem 3.1.
+(ii) Now assume that $m$ is odd. Then $\mu_{2}(A_{m})\otimes H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\simeq\mathbb{Z}/2$. Moreover, from the commutative diagram
+${H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}),\mathbb{Z})}$${\mu_{2}(\mathbb{Z})\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})}$${H_{3}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})}$${\mu_{2}(A_{m})\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}$
+it follows that the bottom map is surjective (note that the upper map is surjective in view of Theorem 3.1, because $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})=0$). Since
 $H_{1}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})\simeq(\mathbb{Z}/2)^{\alpha}\oplus(\mathbb{Z}/3)^{\beta}$
-for some $\alpha$ and $\beta$ (see [31, Corollary 4.4]), from the group structure of $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ one sees that the natural map $\mu_{2}(A_{m})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ should be injective.
-Hence the claim follows from Theorem 2.1.
-∎
-Example 3.2.
-Let $A_{m}:=\mathbb{Z}[\frac{1}{m}]$, where $m=2$ or $3$.
-(i) The product map $\mu_{2}(A_{2})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})$ is trivial: By [23, Corolllary 3.7, Remark 3.8] we have the isomorphism $H_{3}(\mathit{{\rm PSL}}_{2}(A_{2}),\mathbb{Z})\simeq\mathcal{RB}(A_{2})$, where $\mathcal{RB}(A_{2})$ is the refined Bloch group of $A_{2}$ (see the next section).
-By a result of Coronado and Hutchinson, we have the refined Bloch-Wigner exact sequence
-$0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})\rightarrow\mathcal{RB}(A_{2})\rightarrow 0$
-(see [8, Proposition 8.31]). From these two results it follows that the desired product map is trivial.
-(ii) Now we show that the product map $\mu_{2}(A_{3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})$ is not trivial: We know that $H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/12\oplus\mathbb{Z}/4$.
-This follows from the cohomology calculation of $\mathit{{\rm SL}}_{2}(A_{3})$ in [1, page 8] and Lemma 3.3 below.
-Furthermore, by [31, Proposition 4.2], $H^{k}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ is generated by $2$-torsion and $3$-torsion elements for $k\geq 3$.
-Again by Lemma 3.3 it follows that $H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ is generated by $2$-torsion and $3$-torsion elements.
-Now let the above product map is trivial.
-Then from Proposition 3.1 we obtain the exact sequence
-$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0.\end{array}$
-This and the isomorphism $H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/12\oplus\mathbb{Z}/4$ imply that $H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ has an element of order $4$, which is a contradiction.
+for some $\alpha$ and $\beta$ (see [28, Corollary 4.4]), from the group structure of $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ one sees that the natural map
+$\mu_{2}(A_{m})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$
+must be injective. Hence the claim follows from Theorem 3.1. ∎
+Remark 4.2.
+A central extension $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$ is called a stem-extension if the natural map
+$\mathcal{A}=H_{1}(\mathcal{A},\mathbb{Z})\rightarrow H_{1}(\mathcal{G},\mathbb{Z})=\mathcal{G}^{\rm ab}$
+is trivial [13, §4] and it is called a weak stem-extension if the natural map
+$\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})$
+is trivial [13, page 103]. The proof of Theorem 3.1 shows that if $A$ is a ring such that $\mu_{2}(A)=\{\pm 1\}$, then
+$1\rightarrow\mu_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)\rightarrow 1$
+always is a weak stem-extension.
+But by Proposition 4.1 this extension is not a stem-extension for $A=\mathbb{Z}[\frac{1}{m}]$, where $m$ is a square free odd integer.
+This also is true for $A=\mathbb{Z}$ by the paragraph above Proposition 4.1 and $A=\mathbb{Z}_{(2)}$ by Example 5.3.
 We learned the following result from Kevin Hutchinson. It seems that this lemma is widely known, but we could not find any suitable reference for it.
@@
 ∎
-4. A refined Bloch-Wigner exact sequence
-Let $A$ be a commutative ring.
-For any non-negative integer $n$, let $X_{n}(A^{2})$ be the free abelian group generated by the set of all $(n+1)$-tuples $(\langle{\boldsymbol{v}_{0}}\rangle,\dots,\langle{\boldsymbol{v}_{n}}\rangle)$, where every ${\boldsymbol{v}_{i}}\in A^{2}$ is unimodular and for any two distinct vectors ${\boldsymbol{v}_{i}},{\boldsymbol{v}_{j}}$, the matrix $({\boldsymbol{v}_{i}},{\boldsymbol{v}_{j}})$ is invertible.
-We consider $X_{l}(A^{2})$ as a left $\mathit{{\rm SL}}_{2}(A)$-module (resp.
-left $\mathit{{\rm PSL}}_{2}(A)$-module) in a natural way.
-We convert this action to a right action by the definition $m.g:=g^{-1}m$.
-The $l$-th differential operator
-$\partial_{l}:X_{l}(A^{2})\rightarrow X_{l-1}(A^{2}),\ \ l\geq 1,$
-is defined as an alternating sum of face operators which throws away the $i$-th component of generators.
-Let $\partial_{-1}=\epsilon:X_{0}(A^{2})\rightarrow\mathbb{Z}$ be defined by $\sum_{i}n_{i}(\langle{\boldsymbol{v}_{0,i}}\rangle)\mapsto\sum_{i}n_{i}$.
-Hence we have the complex
-$X_{\bullet}(A^{2})\rightarrow\mathbb{Z}:\ \cdots\longrightarrow X_{2}(A^{2})\overset{\partial_{2}}{\longrightarrow}X_{1}(A^{2})\overset{\partial_{1}}{\longrightarrow}X_{0}(A^{2})\overset{\epsilon}{\longrightarrow}\mathbb{Z}\rightarrow 0.$
-Proposition 4.1 (Hutchinson).
-If $A$ is a local ring, then $X_{\bullet}(A^{2})\rightarrow\mathbb{Z}$ is exact in dimension $<|A/\mathfrak{m}_{A}|$.
-Proof.
-See [16, Lemma 3.21]. ∎
-From now on we assume that $A$ is a local ring. Let $Z_{2}(A^{2}):=\ker(\partial_{2})$ and set
-$\mathcal{RP}(A):=H_{0}(\mathit{{\rm SL}}_{2}(A),Z_{2}(A^{2}))=H_{0}(\mathit{{\rm PSL}}_{2}(A),Z_{2}(A^{2})).$
-If the residue field of $A$ has at least 4 elements, then $\mathcal{RP}(A)$ is isomorphic with the quotient of the free $\mathcal{G}_{A}$-module generated by the symbols $[x]$, $x\in\mathcal{W}_{A};=\{a\in A:a(1-a)\in{A^{\times}}\}$, over the subgroup generated by the elements
-$[x]-[y]+\langle x\rangle\bigg{[}\frac{y}{x}\bigg{]}-\langle x^{-1}-1\rangle\Bigg{[}\frac{1-x^{-1}}{1-y^{-1}}\Bigg{]}+\langle 1-x\rangle\Bigg{[}\frac{1-x}{1-y}\Bigg{]},\ \ \ \ \ \ x,y,x/y\in\mathcal{W}_{A}$
-(see [8, §6] and [21, §2]). The inclusion ${\rm inc}:Z_{2}(A^{2})\rightarrow X_{2}(A^{2})$ induces the map
-$\lambda:\mathcal{RP}(A)=Z_{2}(A^{2})_{\mathit{{\rm SL}}_{2}(A)}\overset{\overline{{\rm inc}}}{\longrightarrow}X_{2}(A^{2})_{\mathit{{\rm SL}}_{2}(A)}.$
-The orbits of the action of $\mathit{{\rm SL}}_{2}(A)$ on $X_{2}(A)$ is represented by
-$\langle a\rangle[\ ]:=(\langle{\boldsymbol{e}_{1}}\rangle,\langle{\boldsymbol{e}_{2}}\rangle,\langle{\boldsymbol{e}_{1}}+a{\boldsymbol{e}_{2}}\rangle),\ \ \ \ \langle a\rangle\in\mathcal{G}_{A}.$
-Therefore $X_{2}(A^{2})_{\mathit{{\rm SL}}_{2}(A)}\simeq\mathbb{Z}[\mathcal{G}_{A}]$. The group
-$\mathcal{RP}_{1}(A):=\ker\Big{(}\lambda:\mathcal{RP}(A)\rightarrow\mathbb{Z}[\mathcal{G}_{A}]\Big{)}.$
-is called the scissors congruence group of $A$. We call
-${\rm GW}(A):=H_{0}(\mathit{{\rm SL}}_{2}(A),Z_{1}(A^{2}))$
-the Grothendieck-Witt group of $A$ (see [21, Remark 2.2] for an explanation on the choice of this name). Let
-$\epsilon:=d_{2,0}^{1}:{\rm GW}(A)\rightarrow\mathbb{Z}.$
-The kernel of $\epsilon$ is called the fundamental ideal of $A$ and is denoted by $I(A)$.
-Let $\mathcal{I}_{A}$ be the augmentation ideal of $\mathbb{Z}[\mathcal{G}_{A}]$.
-It is easy to see that $\mathcal{I}_{A}$ is generated by the elements $\langle\!\langle a\rangle\!\rangle:=\langle a\rangle-1$, $a\in{A^{\times}}$ and the natural map $\mathbb{Z}[\mathcal{G}_{A}]\rightarrow{\rm GW}(A)$ is surjective and maps $\mathcal{I}_{A}$ onto $I(A)$.
-If $|A/\mathfrak{m}_{A}|\neq 2$, then
-${\rm GW}(A)\simeq\mathbb{Z}[\mathcal{G}_{A}]/\langle\langle\!\langle a\rangle\!\rangle\langle\!\langle 1-a\rangle\!\rangle:a,1-a\in{A^{\times}}\rangle,$
-$I(A)\simeq\mathcal{I}_{A}/\langle\langle\!\langle a\rangle\!\rangle\langle\!\langle 1-a\rangle\!\rangle:a,1-a\in{A^{\times}}\rangle$
-(see [21, §2]).
-Let $L_{\bullet}(A^{2})$ be the complex $0\rightarrow Z_{1}(A^{2})\overset{{\rm inc}}{\rightarrow}X_{1}(A^{2})\overset{\partial_{1}}{\rightarrow}X_{0}(A^{2})\rightarrow 0$.
-From this we obtain the first quadrant spectral sequences
-$E^{1}_{p.q}=H_{q}(\mathit{{\rm PSL}}_{2}(A),L_{p}(A^{2}))\Longrightarrow H_{p+q}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z}),$
-${E^{\prime}}^{1}_{p.q}=H_{q}(\mathit{{\rm SL}}_{2}(A),L_{p}(A^{2}))\Longrightarrow H_{p+q}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}).$
-These spectral sequences have been studies extensively in [23] and [21]. One can show that
-$E_{0,q}^{1}\simeq H_{q}(\mathit{{\rm PB}}(A),\mathbb{Z}),\ \ \ \ \ \ E_{1,q}^{1}\simeq H_{q}(\mathit{{\rm PT}}(A),\mathbb{Z}),$
-${E^{\prime}}_{0,q}^{1}\simeq H_{q}(B(A),\mathbb{Z}),\ \ \ \ \ \ {E^{\prime}}_{1,q}^{1}\simeq H_{q}(T(A),\mathbb{Z}),$
-where
-$T(A):=\bigg{\{}\begin{pmatrix}a&0\\ 0&a^{-1}\end{pmatrix}:a\in{A^{\times}}\bigg{\}}\ \ \subseteq\ \ B(A):=\bigg{\{}\begin{pmatrix}a&b\\ 0&a^{-1}\end{pmatrix}:a\in{A^{\times}},b\in A\bigg{\}}.$
-and
-$\mathit{{\rm PB}}(A):=B(A)/\mu_{2}(A)I_{2},\ \ \ \ \ \ \mathit{{\rm PT}}(A):=T(A)/\mu_{2}(A)I_{2}.$
-We denote the matrix $\begin{pmatrix}a&0\\ 0&a^{-1}\end{pmatrix}$ by $D(a)$.
-Proposition 4.2 (Hutchinson).
-Let $A$ be a local ring (local domain) such that either $A/\mathfrak{m}_{A}$ is infinite or if $|A/\mathfrak{m}_{A}|=p^{d}$, we have $(p-1)d>2(n+1)$ ($(p-1)d>2n$).
-Then $H_{n}(T(A),\mathbb{Z})\simeq H_{n}(B(A),\mathbb{Z})$ and $H_{n}(\mathit{{\rm PT}}(A),\mathbb{Z})\simeq H_{n}(\mathit{{\rm PB}}(A),\mathbb{Z})$.
-Proof.
-See [21, Proposition 3.8] and [23, Lemma 2.8]. ∎
-From now on we assume that $A$ is a local ring (local domain) such that $A/\mathfrak{m}_{A}$ either is infinite or if it has $p^{d}$ elements, then $(p-1)d>8$ ($(p-1)d>6$).
-Then we have the isomorphisms
-$H_{n}(T(A),\mathbb{Z})\simeq H_{n}(B(A),\mathbb{Z})$
-$H_{n}(\mathit{{\rm PT}}(A),\mathbb{Z})\simeq H_{n}(\mathit{{\rm PB}}(A),\mathbb{Z})$
-for $n\leq 3$.
-It is shown in [23, Lemma 2.1] that $E_{2,1}^{2}\simeq\mathcal{RP}_{1}(A)$ and $E_{0,2}^{2}\simeq\displaystyle\frac{{A^{\times}}\wedge{A^{\times}}}{{A^{\times}}\wedge\mu_{2}(A)}$.
-The kernel of
-$\lambda_{1}:=d_{2,1}^{2}:\mathcal{RP}_{1}(A)\rightarrow\frac{{A^{\times}}\wedge{A^{\times}}}{{A^{\times}}\wedge\mu_{2}(A)}$
-is denoted by $\mathcal{RB}(A)$ and is called the refined Bloch group of $A$ (see [23, Lemma 2.1 and page 8]). Moreover, it is shown in [21, page 11] that the sequence
-[…]
-is exact. By studying the spectral sequence $E_{\bullet,\bullet}^{1}$ we proved the following projective refined Bloch-Wigner exact sequence in [23].
-Theorem 4.3.
-Let $A$ be a local ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu(A)\rightarrow\mu(F)$ is injective.
-Assume that $A/\mathfrak{m}_{A}$ either is infinite or if it has $p^{d}$ elements, then $(p-1)d>8$ (normal-((($(p-1)d>6$ when $A$ is a domain)normal-))).
-If either $-1\in{A^{\times}}^{2}$ or $|\mathcal{G}_{A}|\leq 4$, then we have the exact sequence
-$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(A),\widetilde{\mu}(A))\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mathcal{RB}(A)\rightarrow 0,$
-where $\widetilde{\mu}(A)=\mu(A)/\mu_{2}(A)$.
-Proof.
-See [23, Theorem 3.1, Theorem 3.2, Corollary 3.3] and [21, Proposition 3.8]. ∎
-Now we have the following generalization of Hutchinson’s refined Bloch-Wigner exact sequence.
-Theorem 4.4.
-Let $A$ be a local ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu(A)\rightarrow\mu(F)$ is injective.
-Assume that $A/\mathfrak{m}_{A}$ either is infinite or if it has $p^{d}$ elements, then $(p-1)d>8$ (normal-((($(p-1)d>6$ when $A$ is a domain)normal-))).
-If either $-1\in{A^{\times}}^{2}$ or $|\mathcal{G}_{A}|\leq 4$, then we have the exact sequence
-$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)}\rightarrow\mathcal{RB}(A)\rightarrow 0$
-Proof.
-The composite
-$\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\twoheadrightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\twoheadrightarrow\mathcal{RB}(A)$
-is surjective. Let $\mathcal{K}$ be the kernel of this map. By Theorem 4.3 we have the commutative diagram with exact rows
-[…]
-From this diagram and Corollary 2.4 we obtain the exact sequence
-$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\mathcal{K}\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(A),\widetilde{\mu}(A))\rightarrow 0$
-This clearly implies that $\mathcal{K}\simeq{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}$.
-Now we study the product map $\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$.
-Since $A/\mathfrak{m}_{A}$ has at least $4$ elements, we have $A_{A^{\times}}=H_{0}({A^{\times}},A)=0$, where the action of ${A^{\times}}$ on $A$ is given by $a.x:=a^{2}x$.
-First note that ${E^{\prime}}_{0,1}^{1}\simeq\mathcal{G}_{A}\oplus A_{A^{\times}}=\mathcal{G}_{A}$.
-The map ${d^{\prime}}_{2,0}^{2}:I(A)\rightarrow\mathcal{G}_{A}$ is given by $\langle\!\langle a\rangle\!\rangle\mapsto\langle a\rangle$ [8, Lemma 4.7].
-Thus ${E^{\prime}}_{2,0}^{\infty}\simeq{E^{\prime}}_{2,0}^{2}\simeq I^{2}(A)$.
-Now from an easy analysis of the spectral sequence ${E^{\prime}}_{\bullet,\bullet}^{1}$ we obtain the exact sequence
-$H_{2}(B(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A)\rightarrow 0.$
-By Proposition 4.2, $H_{2}(T(A),\mathbb{Z})\simeq H_{2}(B(A),\mathbb{Z})$. The composite
-$\mu_{2}(A)\otimes(T(A)\wedge T(A))\simeq\mu_{2}(A)\otimes H_{2}(T(A),\mathbb{Z})\rightarrow\mu_{2}(A)\otimes H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
-$\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
-is given by $(-1)\otimes(D(a)\wedge D(b))\mapsto{\boldsymbol{c}}(D(-1),D(a),D(b)).$ If $i^{2}=-1$, then
-${\boldsymbol{c}}(D(-1),D(a),D(b)=2{\boldsymbol{c}}(D(i),D(a),D(b))=0.$
-Let $|\mathcal{G}_{A}|\leq 4$. If $-1\in{A^{\times}}^{2}$, we proceed as in above. If $-1\notin{A^{\times}}^{2}$, then $\mathcal{G}_{A}=\{1,-1,u,-u\}$ for some $u\in{A^{\times}}$. Thus
-${\boldsymbol{c}}(D(-1),D(a),D(b))={\boldsymbol{c}}(D(-1),D(-1),D(u))=0.$
-In any case, the above composite is trivial and thus
-$\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}=\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)}.$
-This completes the proof of the theorem. ∎
+Example 4.4.
+Let $A_{m}:=\mathbb{Z}[\frac{1}{m}]$, where $m=2$ or $3$.
+(i) The product map $\mu_{2}(A_{2})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})$ is trivial: By [20, Corolllary 3.7, Remark 3.8] we have the isomorphism
+$H_{3}(\mathit{{\rm PSL}}_{2}(A_{2}),\mathbb{Z})\simeq\mathcal{RB}(A_{2}),$
+where $\mathcal{RB}(A_{2})$ is the refined Bloch group of $A_{2}$ (see [8, §4] for the definition of $\mathcal{RB}(A_{2})$).
+By a result of Coronado and Hutchinson, we have the refined Bloch-Wigner exact sequence
+$0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})\rightarrow\mathcal{RB}(A_{2})\rightarrow 0$
+(see [8, Proposition 8.31]). From these two results and Proposition 4.1, it follows that the desired product map is trivial.
+(ii) Now we show that the product map $\mu_{2}(A_{3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})$ is not trivial.
+We know that
+$H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/12\oplus\mathbb{Z}/4.$
+This follows from the cohomology calculation of $\mathit{{\rm SL}}_{2}(A_{3})$ in [1, page 8] and Lemma 4.3.
+Furthermore, by [28, Proposition 4.2], $H^{k}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ is generated by elements of orders $2$ and $3$ for $k\geq 3$.
+Again by Lemma 4.3 it follows that $H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ is generated by elements of orders $2$ and $3$.
+Now let the above product map be trivial.
+Then by Proposition 4.1 the sequence
+$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0\end{array}$
+must be exact.
+This and the isomorphism $H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/12\oplus\mathbb{Z}/4$ imply that $H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ has an element of order $4$, which is a contradiction.
+Observe that $H_{2}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}$ and thus $\mu_{2}(A_{3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/2$.
+5. The low dimensional homology groups of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$ of local rings
+Let $A$ be a local ring with maximal ideal $\mathfrak{m}_{A}$. It is known that
+$H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq\begin{cases}A/\mathfrak{m}_{A}^{2}&\text{if $|A/\mathfrak{m}_{A}|=2$}\\ A/\mathfrak{m}_{A}&\text{if $|A/\mathfrak{m}_{A}|=3$.}\\ 0&\text{if $|A/\mathfrak{m}_{A}|\geq 4$}\\ \end{cases}$
+(see [19, Proposition 4.1]). Now let $|A/\mathfrak{m}_{A}|\neq 2$. Since $\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})=0$, from Theorem 1.1 we obtain the exact sequence
+$0\rightarrow H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mu_{2}(A)\rightarrow 0$
+Since
+$\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})=0,\ \ \ \ {{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(A),H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}))=0,$
+by Theorem 3.1 we obtain the following result.
+Proposition 5.1.
+Let $A$ be a local ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu_{2}(A)\simeq\mu_{2}(F)$. If $|A/\mathfrak{m}_{A}|\neq 2$, then the sequence
+$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0$
+is exact.
 Remark 5.2.
-We actually can calculate the map
-$\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
-which appears in the above theorem.
-Let $A$ be a local domain such that ${\rm char}(A)\neq 2$.
-Let $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in I^{2}(A)$.
-Then by some tedious calculation one can show that under the surjective map $H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A)$, the cycle class of
-$[a|b]+R_{ab}-R_{a}-R_{b}+R_{1}\in B_{2}(\mathit{{\rm SL}}_{2}(A))_{\mathit{{\rm SL}}_{2}(A)}$
-maps to $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in I^{2}(A)$, where for any $z\in{A^{\times}}$,
-$R_{z}$ $=[w|g^{-1}_{z}]-[w|h^{-1}_{z}]-[h^{-1}_{z}D(z)^{-1}|wh^{-1}_{z}]+[h_{z^{-1}}|w^{-1}]-[h^{-1}_{z}|D(z)^{-1}]$
-$+[D(z)h_{z}|D(z)^{-1}]-[D(z)h_{z}|h^{-1}_{z}]+2[h^{-1}_{z}|h^{-1}_{z}]+[h^{-2}_{z}|h^{-2}_{z}]$
-$+[D(2)h^{-1}_{z}D(z)^{-1}|D(2)]-[D(2)|h^{-1}_{z}],$
-with $g_{z}=\left(\!\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!z\end{array}\!\!\right)$ and $h_{z}=\left(\!\!\begin{array}[]{cc}\!\!1&\!\!z^{-1}\\ \!\!0&\!\!1\end{array}\!\!\right)$.
-For another cycle that its class maps to $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in I^{2}(A)$ see [15, §7].
-Now $(-1)\otimes\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle$ maps in $H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ to the class of the cycle which is obtained from the shuffle product of $[-1]\in B_{1}(\mu_{2}(A))_{\mu_{2}(A)}$ with the cycle $[a|b]+R_{ab}-R_{a}-R_{b}+R_{1}$.
+The result of the above proposition is valid for any domain $A$ such that $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is torsion of odd exponent.
+For example let $\mathcal{O}_{-3}$ be the ring of algebraic integers of the imaginary quadratic field $\mathbb{Q}(\sqrt{-3})$.
+The homology groups of $\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3})$ and $\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3})$ have been calculated in [2, Theorem, page 376] and [25, Theorem 5.7].
+We list these homology groups for $1\leq n\leq 3$:
+$H_{n}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\simeq\begin{cases}\mathbb{Z}/3&\text{if $n=1$}\\ \mathbb{Z}/4&\text{if $n=2,$}\\ \mathbb{Z}/24\times\mathbb{Z}/6&\text{if $n=3$}\end{cases}$
+$H_{n}(\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\simeq\begin{cases}\mathbb{Z}/3&\text{if $n=1$}\\ \mathbb{Z}/4\times\mathbb{Z}/2&\text{if $n=2.$}\\ \mathbb{Z}/6\times\mathbb{Z}/3&\text{if $n=3$}\end{cases}$
+From Theorem 3.1 we have the exact sequence
+$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})}{\mu_{2}(\mathcal{O}_{-3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\rightarrow 0.\end{array}$
+From the sizes of the third homology groups of $\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3})$ and $\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3})$ it follows that the product map
+$\mu_{2}(\mathcal{O}_{-3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3}),\mathbb{Z})$
+is non-trivial.
+Example 5.3.
+Let $A$ be a local domain such that $A/\mathfrak{m}_{A}\simeq\mathbb{F}_{2}$. Then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq A/\mathfrak{m}_{A}^{2}$. By a careful analysis of the composition
+$\mu_{2}(A)\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\overset{\simeq}{\longrightarrow}A/\mathfrak{m}_{A}^{2},$
+one sees that $-1\in\mu_{2}(A)$ maps to $-\overline{6}\in A/\mathfrak{m}_{A}^{2}$.
+Since $2\in\mathfrak{m}_{A}$, we have $-\overline{6}=\overline{2}$ in $A/\mathfrak{m}_{A}^{2}$.
+Thus if $2\in\mathfrak{m}_{A}^{2}$, then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq H_{1}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})$.
+Otherwise we have the exact sequence
+$1\rightarrow\mu_{2}(A)\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0$
+Observe that
+$\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq\mu_{2}(A)\otimes_{\mathbb{Z}}A/\mathfrak{m}_{A}^{2}\simeq A/(2A+\mathfrak{m}_{A}^{2}).$
+If $A=\mathbb{Z}_{(2)}=\{a/b\in\mathbb{Q}:a,b\in\mathbb{Z},2\nmid b\}$ (or $A=\mathbb{Z}_{2}$, the ring of $2$-adic integers), then $2\notin\mathfrak{m}_{A}^{2}$ (by the Nakayama lemma).
+Thus the sequence
+$1\rightarrow\mu_{2}(\mathbb{Z}_{(2)})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow 0$
+is exact. Note that $H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq\mathbb{Z}/4$ [19, Example 4.2]. With an argument similar to the proof of Proposition 4.1(ii), one can show that
+${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(\mathbb{Z}_{(2)}),H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z}))\rightarrow\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}{\mu_{2}(\mathbb{Z}_{(2)})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}$
+is trivial and
+$H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow\mu_{2}(\mathbb{Z}_{(2)})\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq\mathbb{Z}/2$
+is surjective. Thus
+$H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq H_{2}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})$
+and the sequence
+$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}{\mu_{2}(\mathbb{Z}_{(2)})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0\end{array}$
+is exact. We have similar results for $\mathbb{Z}_{2}$. Hence Proposition 5.1 is not necessarily true when $|A/\mathfrak{m}_{A}|=2$.
 Proposition 5.4.
-If $F$ is a real closed field, then the sequence
-$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{RB}(F)\rightarrow 0$
-is exact.
-Proof.
-It is well-know that for any real closed field $F$, $I(F)\simeq\mathbb{Z}$ which is generated by $\langle\!\langle-1\rangle\!\rangle$ and $I^{2}(F)=2I(F)\simeq 2\mathbb{Z}$.
-Consider the exact sequence
-[…]
-(discussed before Theorem 4.3). Note that
-$I(F)\otimes_{\mathbb{Z}}\mu_{2}(F)\simeq I(F)/2I(F)\simeq I(F)/I^{2}(F)\simeq\mathcal{G}_{F}\simeq\mathbb{Z}/2.$
-Applying the Snake lemma to the commutative diagram
-[…]
-we obtain the exact sequence
-[…]
-We show that ${E^{\prime}}_{1,2}^{2}=0$. Consider the differential
-${d^{\prime}}_{2,2}^{1}:H_{2}(\mathit{{\rm SL}}_{2}(F),Z_{1}(F^{2}))\rightarrow H_{2}(T(A),\mathbb{Z})\simeq F^{\times}\wedge F^{\times}=F^{>0}\wedge F^{>0}$
-(note that $F=\mu_{2}(F)\oplus F^{>0}$ and $F^{>0}=(F^{>0})^{2}$). For any $a,b\in F^{>0}$, let
-$Y_{a,b}:=\overline{([D(a)|D(b)]-[D(b)|D(a]))\otimes\bigg{(}(\langle{\boldsymbol{e}_{1}}\rangle,\langle{\boldsymbol{e}_{2}}\rangle)+(\langle{\boldsymbol{e}_{2}}\rangle,\langle{\boldsymbol{e}_{1}}\rangle)\bigg{)}}\in H_{2}(\mathit{{\rm SL}}_{2}(F),Z_{1}(F^{2})).$
-It is straightforward to check that
-${d^{\prime}}_{2,2}^{1}(Y_{a,b})=2(a\wedge b).$
-Since $F^{>0}=(F^{>0})^{2}$, ${d^{\prime}}_{2,2}^{1}$ is surjective and thus ${E^{\prime}}_{1,2}^{2}=0$.
-Moreover, it is not difficult to show that ${E^{\prime}}_{0,3}^{2}\simeq(\bigwedge_{\mathbb{Z}}^{3}F^{\times})/2\oplus{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))={{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))$ (see [23, Proposition 2.7] and the paragraph after that).
-Now by an easy analysis of the spectral sequence ${E^{\prime}}_{\bullet,\bullet}^{1}$, we obtain the exact sequence
-$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow{E^{\prime}}^{\infty}_{2,1}\rightarrow 0$
-From the commutative diagram with exact rows
-[…]
-it follows that the middle vertical map has a kernel with at most four elements. But we have seen that the composite map
-$\mathbb{Z}/4\simeq H_{3}(\langle w\rangle,\mathbb{Z})\hookrightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{RB}(F)$
-is trivial (see the proof of Theorem 2.1).
-This shows that the kernel of the surjective map $H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\twoheadrightarrow\mathcal{RB}(F)$ has exactly four elements.
-Hence we are done.
-∎
-Corollary 4.7.
-If $F$ is a real closed field, then the sequence
-$0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\rightarrow 0$
+Let $F$ be either a finite field, a quadratically closed field, or a real closed field. Then the sequence
+$0\rightarrow\mu_{2}(F)^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\rightarrow 0$
 is exact. In particular, the product map
@@
 Proof.
-This follows from Proposition 4.6, Theorem 4.3 and Corollary 2.4. ∎
+(i) Let $F=\mathbb{F}_{q}$ be finite. For fields of characteristic $2$ the above exact sequence is trivial. Thus we may assume that $q$ is odd. We have seen that
+$H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})\simeq\begin{cases}0&\text{if $q\neq 3$}\\ \mathbb{Z}/3&\text{if $q=3$.}\end{cases}$
+Moreover, it is well-known that
+$H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})\simeq\begin{cases}0&\text{if $q\neq 9$}\\ \mathbb{Z}/3&\text{if $q=9$}\end{cases}$
+(see [24, page 280]). These facts together with Theorem 3.1 give us the desired exact sequence.
+(ii) Let $F$ be quadratically closed. By Proposition 2.2, $H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})$ is uniquely 2-divisible. Thus by Proposition 5.1 we obtain the desired exact sequence.
+(iii) Let $F$ be a real closed field. By Proposition 2.4 we have the Bloch-Wigner exact sequence
+$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{B}(F)\rightarrow 0$
+From [20, Corollary 3.5] and [7, Corollary 4.7] we obtain the isomorphism
+$H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\simeq\mathcal{B}(F).$
+Our claim follows from these two facts. Thus we presented a proof of this in above proposition. ∎
+Remark 5.5.
+It is mentioned in [11, page 230] that the sequence
+$0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{R}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{R}),\mathbb{Z})\rightarrow 0$
+is exact and for a proof the authors referred to [10, App. A]. We couldn’t really verify this. So we gave a proof of this in above proposition.